### v1

# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

Abstract
Spatio-temporal forecasting is crucial in many domains, such as transportation, meteorology, and energy. However, real-world scenarios frequently present challenges such as signal anomalies, noise, and distributional shifts. Existing solutions primarily enhance robustness by modifying network architectures or training procedures. Nevertheless, these approaches are computationally intensive and resource-demanding, especially for large-scale applications. In this paper, we explore a novel test-time computing paradigm, namely learning with calibration, ST-TTC, for spatio-temporal forecasting. Through learning with calibration, we aim to capture periodic structural biases arising from non-stationarity during the testing phase and perform real-time bias correction on predictions to improve accuracy. Specifically, we first introduce a spectral-domain calibrator with phase-amplitude modulation to mitigate periodic shift and then propose a flash updating mechanism with a streaming memory queue for efficient test-time computation. ST-TTC effectively bypasses complex training-stage techniques, offering an efficient and generalizable paradigm. Extensive experiments on real-world datasets demonstrate the effectiveness, universality, flexibility and efficiency of our proposed method.
1 Introduction
Spatio-temporal forecasting (STF) aims to predict the future state of dynamic systems from historical spatio-temporal observations and underpins many real-world applications, such as traffic flow forecasting guo2021learning , air quality forecasting nguyen2023climax , and energy consumption forecasting wang2019review . Although spatio-temporal neural networks shao2024exploring ; jin2023spatio ; jin2024survey , which couple spatial neural operators with temporal neural operators, have achieved remarkable progress on these tasks, their deployment in practical environments remains fraught with challenges. These observations, typically collected by sensors, are frequently corrupted by noise, outliers (e.g., spikes or dropouts due to hardware failure) xu2024spatiotemporal , and more commonly, non-stationary distribution shifts arising from sensor aging and seasonal patterns wang2024evaluating .
To enhance generalization and performance, prior work has focused primarily on out-of-distribution (OOD) learning for ST data during the training phase: designing architectures that resist perturbations ji2025seeing ; ma2024stop ; wang2024stone ; zhou2024coms2t ; zhou2023maintaining ; xia2023deciphering , augmenting training data with noise or adversarial examples bai2025alleviating ; zhang2023automated ; liu2022practical , and introducing specialized loss functions or regularizers li2024flashst ; zhang2023spatial to counteract distribution drift. However, these methods share fundamental limitations: they assume that the training data sufficiently captures all future target domain invariance—a premise that is rarely valid in real-world settings. Concurrently, an emerging paradigm of continual fine-tuning chen2025eac ; yi2024get ; wang2024towards ; lee2024continual ; wang2023pattern ; wang2023knowledge ; kieu2024team ; chen2021trafficstream has become popular in spatio-temporal learning by continuously tuning the model to adapt to dynamic changes. Though promising, it still divides the target domain into multiple periods of training and testing and relies on period-specific training data to optimize model, thereby failing in data-scarse scenarios.
Table 1: Formal comparison of different spatio-temporal learning paradigms for generalization from the perspective of data and learning. $s$ denotes the source domain, $t$ denotes the target domain, $x$ and $y$ denote the samples and labels sampled from $\mathbf{X}$ and $\mathbf{Y}$, respectively. OOD learning expects inputs sampled from any environment $e^{*}\sim\mathcal{E}$ to be valid, while others are only optimized for the current training or test environment $e$. In particular, continual fine-tuning divides the target domain into multiple stages and optimizes for a specific stage $\tau$ environment $e^{\tau}$.  means not involved.
Setting Example Works Data Perspective Learning Perspective
Source Target Train-Time Test-Time
OOD Learning STONE wang2024stone , CaST xia2023deciphering $\langle\mathbf{X}^{s},\mathbf{Y}^{s}\rangle$ $\min_{f_{\theta}}\;\max_{e^{*}\in\mathcal{E}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{s},\mathbf{Y}^{s}\mid e^{*})}\bigl{[}L(f_{\theta}(x),y)\bigr{]}$
Continual Fine-Tuning EAC chen2025eac , TrafficStream chen2021trafficstream $\langle\mathbf{X}^{t},\mathbf{Y}^{t}\rangle$ $\min_{f_{{\theta}^{\tau}}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t},\mathbf{Y}^{t}\mid e^{\tau})}\bigl{[}L(f_{{\theta}^{\tau}}(x),y)\bigr{]}$
Test-Time Training TTT-ST chen2024test $\mathbf{X}^{s}$ $\mathbf{X}^{t}$
$\min_{f_{\theta}}\,\mathbb{E}_{(\tilde{x},x)\sim P(\mathbf{X}^{s}\mid e)}\bigl{[}L(f_{\theta}(\tilde{x}),x)\bigr{]}$
$\min_{f_{\theta}}\,\mathbb{E}_{(\tilde{x},x)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L(f_{\theta}(\tilde{x}),x)\bigr{]}$
Online Continual Learning DOST wang2024distribution $\mathbf{X}^{t}$ $\min_{f_{\theta(\delta)}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L(f_{\theta(\delta)}(x),y)\bigr{]}$
Test-Time Computing ST-TTC (Our) $\mathbf{X}^{t}$ $\min_{g_{\theta}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L\bigl{(}g_{\theta}(f_{\theta}(x)),y\bigr{)}\bigr{]}$
Figure 1: Conceptual visualization comparison of different spatio-temporal learning paradigms under test environment. (a) Test-Time Training requires the use of additional pretext tasks in the training and test phases to optimize the self-supervision head or the overall model parameters $f_{\theta}$. (b) Online Continual Learning, by optimizing some internal parameters $f_{\theta(\delta)}$ of the model, requires additional modifications to the internal architecture of the network. Our (c) Test-Time Computing method only requires a lightweight calibrator $g_{\theta}$, which is a seamless and lightweight plug-and-play module.
Recently, leveraging test-time information has attracted widespread attention for its ability to significantly improve language model performance on complex reasoning tasks snell2025scaling ; akyurek2024surprising . In computer vision, this concept has already been extensively developed: test-time training (TTT) was first introduced by sun2020test , which defines an auxiliary self-supervised task applied to both training and test samples to better balance bias and variance gandelsman2022test . A similar idea was adapted to spatio-temporal forecasting in TTT-ST chen2024test . Unlike language and vision settings—where obtaining ground-truth labels for test samples at inference time is nearly impossible—STF benefits from label autocorrelation cressie2011statistics : each observation strongly depends on its predecessor, and training instances are constructed from sliding windows, which provide access to historical samples and their true labels. Moreover, this property makes STF also require timeliness shao2024stemo , that is, the additional computing time during inference must be less than the window-stride interval. A recent STF method, DOST wang2024distribution , explores online continual learning, which initially explored this direction. It uses historical test sample labels to dynamically adapt the modified model architecture. Though promising, these approaches typically involve complex self-supervised tasks or structural adaptations and still fall short of the timeliness demands of STF.
To address this gap, we propose Test-Time Computing of Spatio-Temporal Forecasting (ST-TTC), an attractive complementary paradigm. ST-TTC achieves learning with calibration by iteratively leveraging available test information during inference, enabling seamless integration with diverse models. This adapts the model to evolving spatio-temporal patterns, thereby calibrating predictions. Our principal insight is that performance degradation during test time is primarily driven by non-stationary distributional shifts stemming from progressive periodic biases. Therefore, we propose a spectral domain calibrator. This involves appending a lightweight module, operating in the frequency domain, subsequent to the backbone network. This module calibrates biases by learning minor, node-specific amplitude and phase correction factors. Furthermore, a flash gradient updating mechanism with a streaming memory queue, ensures universal, rapid, and resource-efficient test-time computing. Table 1 provides a formal comparison of our method against existing learning paradigms, and Figure 1 offers a conceptual visualization of learning with test domain. In summary, our contributions are:
  • We propose a novel test-time computing paradigm of spatio-temporal forecasting, termed ST-TTC .
  • We systematically explore the goals and means of achieving this paradigm. Concretely, we introduce a spectral domain calibrator with phase-amplitude modulation to mitigate periodic shift and present a flash updating mechanism with a streaming memory queue for efficient test-time computation.
  • Experimental results on real-world spatio-temporal datasets in different fields, scenarios, and learning paradigms demonstrate the effectiveness and universality of ST-TTC .
2 Related Work
Spatio-Temporal Forecasting. Spatio-temporal sequences can be regarded as spatially extended multivariate time series. Although one can trivially apply multivariate forecasting methods nie2023time ; zeng2023transformers ; biller2003modeling ; box1970distribution independently at each location, such decoupling of spatial and temporal dependencies invariably yields suboptimal results shao2024exploring . Classical spatio-temporal forecasting method instead relies on shallow models or spatio-temporal kernels, including feature-based methods ohashi2012wind ; zheng2015forecasting , state space models cliff1975space ; pfeifer1980starima ; bahadori2014fast , and Gaussian process models flaxman2015machine ; senanayake2016predicting . Unfortunately, the overall nonlinearity of these models is limited, and the high complexity of computation and storage further hinders the availability of massive training instances shi2018machine . In recent years, spatio-temporal neural networks kumar2024spatio ; jin2024survey ; jin2023spatio have been widely adopted to learn the complex dynamics of such systems. Early work concentrated on devising neural operators to extract spatial or temporal correlation yu2018spatio ; wu2019graph ; shao2022spatial ; liu2023spatio ; fang2025efficient and on designing fusion architectures to integrate them li2018diffusion ; guo2019attention ; oreshkin2021fc ; cini2023taming ; deng2024disentangling . More recent efforts have explored domain-invariant representation learning ma2024stop ; wang2024stone ; xia2023deciphering and continual model adaptation chen2025eac ; wang2023pattern ; chen2021trafficstream to better accommodate unseen environmental shifts. However, these methods still depend exclusively on offline training data and thus cannot deliver truly timely and effective adaptation in real settings.
Test-Time Computing. Test-time computation is inspired by the human cognition kahneman2011thinking , in which additional computational effort is allocated during inference to improve task performance. This insight has recently driven considerable interest in the nature language process community, fueled by the success of reasoning-augmented language models (e.g., o1 jaech2024openai and r1 guo2025deepseek ) that activate and adapt internal computations at test time via supervised fine-tuning or reinforcement learning (RL) zhang2025and . While the generalization properties of RL-based adaptation remain debated yue2025does , the notion of supervised learning on unlabeled test data dates back to “transductive learning” Gammerman98lbyt in the 1990s and has demonstrated empirical benefits vapnik1999nature ; bottou1992local . In the computer vision domain, this idea was formalized as Test-Time Training sun2020test , which attaches an auxiliary self-supervised head to enable online adaptation to each test instance—a paradigm subsequently generalized as test-time adaptation wangtent ; jangtest ; wang2022continual ; liang2025comprehensive . However, spatio-temporal forecasting has seen limited exploration of such techniques. TTT-ST chen2024test applies TTT-style auxiliary objectives during training and continues to update at inference, and DOST wang2024distribution further incorporates dynamic learning mechanisms within modified model architectures for test-time updates. In addition, some methods zhang2023test ; guo2024online are conceptually close to ours, such as CompFormer zhang2023test , which proposes a test-time compensated representation learning framework, but still requires access to additional training data. Notably, we formalize the test-time computing of spatio-temporal forecasting, and propose a unified learning-with-calibration framework that is general, lightweight, efficient, and effective for STF at test-time.
For more related work, we provide a more detailed introduction in Appendix A.
3 Preliminaries
Problem Definition. Let $x\in\mathbb{R}^{T\times C}$ denote the multivariate time series recorded at each location sensor, capturing the dynamic observations of $C$ measured features in $T$ consecutive time steps. Stacking these sequences for all $N$ locations yields the spatio-temporal tensor $X\in\mathbb{R}^{N\times T\times C}$. Given historical observations $X^{h}\in\mathbb{R}^{N\times T^{h}\times C}$ (and an optional spatial correlation graph $\mathcal{G}$ representing the spatial relationships of $N$ locations), spatio-temporal forecasting aims to learn a mapping $f_{\theta}:(X^{h},\mathcal{G})\longmapsto X^{f}\in\mathbb{R}^{N\times T^{f}\times C}$, where $X^{f}$ is the signal for the next $T^{f}$ time steps. In practice, according to shao2024exploring ; li2018diffusion , the feature to be predicted is usually only the target variable.
Scenario Definition. In deep learning systems, batch-based testing is typically employed to exploit parallelism. In real-world deployment, however, predictions must be produced for each incoming time-step sample—i.e., with batch size $B$ set to 1. At time index $t$, once the new sliding-window input $X_{t}\in\mathbb{R}^{N\times T^{h}\times C}$ arrives, the true labels for all test samples before time index $t-T^{h}-T^{f}+1$ become available. Thus, test-time computing of spatio-temporal forecasting can leverage this accumulated historical information to enhance the accuracy of the current prediction, while ensuring that any additional computation latency remains below a threshold defined by the sliding-window stride.
4 Methodology
Our test-time computing framework of spatio-temporal forecasting (ST-TTC) integrates two synergistic components: 1) a spectral domain calibrator with phase-amplitude modulation; and 2) a flash gradient update mechanism with streaming memory queue. In this section, we introduce these two key components, respectively, from the perspective of what is computed and how it is computed.
4.1 What to Compute? Spectral Domain Calibrator with Phase-Amplitude Modulation
Motivation. Spatio-temporal data, such as traffic flow and air quality, often exhibit periodic patterns (e.g., daily or weekly cycles). However, in real-world deployments, these patterns are not stationary; they are dynamically influenced by various internal and external factors wang2024robust . Such influences lead to non-stationarities manifesting as fluctuations in amplitude (e.g., increased or decreased traffic peaks due to seasonal changes) or phase shifts (e.g., peak hours are advanced or delayed due to traffic congestion). Pre-training models typically fit fixed periodic patterns during training, which makes them vulnerable to performance degradation under such persistent dynamic changes during inference wang2024evaluating . Therefore, we argue that the goal of test-time computation is: how to design an effective calibrator that can efficiently capture such gradual systematic bias from the pattern to correct the prediction errors caused by non-stationarity, while avoiding overfitting to random noise?
Key Challenges. While correction in the time domain is possible zhang2023test ; guo2024online , it often requires extensive parameterization, leading to increased model complexity and limited ability to capture evolving periodic structures. Moreover, the coupled structural and branching modules wang2024distribution ; chen2024test are prone to overfitting the random noise in the spatio-temporal evolution. To address this, we propose calibration in the spectral domain, where periodic variations are more transparently expressed as changes in the amplitude and phase of specific frequency components. Spectral correction offers a potentially more direct and robust solution. However, this introduces two main challenges: ❶ the degree of non-stationarity varies across spatial nodes; and ❷ full-spectrum parameterization is computationally expensive. The core problem thus becomes how to design a lightweight, spatial-aware calibrator.
Implementation Details. To this end, we formally introduce the spectral domain calibrator (SD-Calibrator), which is a lightweight plug-and-play module that performs spectral domain calibration on the time domain prediction results of the pre-trained model, aiming to achieve efficient test-time computation for spatio-temporal forecasting. Specifically, it can be divided into three steps:
  • Spatial-aware Decomposition. To ensure spatial awareness, we apply a real-to-complex fast Fourier transform (rFFT) along the time dimension of the backbone model’s prediction $\hat{y}\in\mathbb{R}^{B\times N\times T}$, separately for each spatial node. This yields the frequency spectrum: $Y_{f}=\mathrm{rFFT}(\hat{y})\in\mathbb{C}^{B\times N\times M}$, where $M=\tfrac{T}{2}+1$ is the number of unique frequency bins for real-valued signals. Then, we decompose $Y_{f}$ into its amplitude $A=|Y_{f}|\in\mathbb{R}^{B\times N\times M}$ and phase $P=\angle Y_{f}\in\mathbb{R}^{B\times N\times M}$.
  • Group-wise Modulation. To ensure lightweight and balanced spectrum expression, we divide the $M$ frequency bins into $G$ contiguous groups of size $\lfloor M/G\rfloor$, and learn per-group, per-node amplitude and phase offsets $\lambda^{\alpha}\in\mathbb{R}^{G\times N\times 1},\lambda^{\phi}\in\mathbb{R}^{G\times N\times 1}$. For each group $g\in\{1,\dots,G\}$, we apply $A_{g}^{{}^{\prime}}=A_{g}\odot(1+\lambda^{\alpha}_{g}\bigr{)},P_{g}^{{}^{\prime}}=P_{g}+\lambda^{\phi}_{g},$ and reconstruct the spectrum as $Y_{f}^{{}^{\prime}}=\bigcup_{g=1}^{G}A^{{}^{\prime}}_{g}\odot e^{(j\,P^{{}^{\prime}}_{g})}$.
  • Inverse Transform. Finally, the calibrated time-domain signal is obtained by Inverser rFFT $\hat{y}_{cal}=\mathrm{irFFT}\bigl{(}Y_{f})\;\in\;\mathbb{R}^{B\times N\times T},$ along the frequency dimension.
For clarity, we provide a Algorithm workflow 1 and Pytorch-Style Pseudocode 2 in Appendix C.1.
Complexity Analysis. The full-spectrum parameterization learns independent amplitude and phase offsets for each of the $M=T/2+1$ frequency bins and $N$ nodes, totaling $2NM$ parameters. In contrast, our $G$-group design learns only $2NG$ parameters. Since $G$ is a constant and $M$ grows linearly with $T$, $G\ll M$ is usually the case. For large-scale long-term scenario, this significantly reduces memory footprint and gradient update cost while retaining interpretable per-band calibration.
Theoretical Analysis. We also provide a theoretical bound on the output perturbation induced by the SD-Calibrator, ensuring controlled deviation from the original prediction to prevent overfitting (Please refer to Theorem 1 and the proof in Appendix B).
4.2 How to Compute? Flash Gradient Update with Streaming Memory Queue
Motivation. The SD-Calibrator provides an effective mechanism for output correction. To accommodate the dynamic nature of spatio-temporal data, its parameters ($\lambda^{\alpha},\lambda^{\phi}$) must be continuously updated during inference. Fortunately, as we discussed above, due to the streaming nature of spatio-temporal data, unlike Visual and textual tasks, we have access to the true labels of historical samples. However, simply accumulating all historical data for updates is not feasible due to the increasing computational load and memory usage. Therefore, we argue that the key to test-time computation is: How to design an efficient data selection and learning mechanism that leverages appropriate historical information to tuning the SD-Calibrator without incurring a lot of computational overhead?
Key Challenges. Although retrieving similar sequences from historical training databases can partially compensate for prediction errors zhang2023test , this assumption is unrealistic, as only test-time information is available in our scenario. Moreover, selectively storing historical test samples via memory bank primarily serves to mitigate catastrophic forgetting in the backbone model wang2024distribution , which misaligns with the learning objective of our SD-calibrator. To address this, we propose freezing the backbone and updating only the calibrator using recent test samples for efficient test-time computing. However, this strategy introduces two critical challenges: ❶ recent studies lau2025fast have shown that real-time updates may cause information leakage; and ❷ excessive updates can lead to overfitting of the calibration parameters and increased computational burden. The core problem thus becomes how to design a efficient calibration parameter learning mechanism without information leakage.
Implementation Details. To address these challenges, we introduce the flash gradient update strategy coupled with a streaming memory queue. The process is as follows:
  • Streaming Memory Queue. We maintain a first-in, first-out (FIFO) queue, denoted as $\mathcal{Q}$, with a maximum size equal to the prediction horizon $T^{f}$. For each incoming test instance $t$, after making a prediction, we store the input-label pair $(X_{t},Y_{t})$ into $\mathcal{Q}$ (Here is for engineering convenience. In real deployment, data points can be merged at each step to form the true label). Once $\mathcal{Q}$ is full, for every new test sample $(X_{n},Y_{n})$ added, the oldest sample pair $(X_{o},Y_{o})$ is dequeued. This dequeued sample $(X_{o},Y_{o})$ is then used for the gradient update, thus avoiding the information leakage.
  • Flash Gradient Update. Once we have $(X_{o},Y_{o})$, we first obtain the backbone model’s prediction for the historical input: $\hat{Y}_{o}^{b}=f_{\theta}(X_{o})$ (note: the backbone model weights $f_{\theta}$ are frozen). Then, the SD-Calibrator $g_{\theta}$ processes this prediction: $\hat{Y}_{o}^{cal}=g_{\theta}(\hat{Y}_{o}^{b})$. The loss function between the calibrated prediction $\hat{Y}_{o}^{cal}$ and the true historical label $Y_{o}$ is calculated, and only a single gradient descent step is performed to update the parameters of the SD-Calibrator: $\lambda\leftarrow\lambda-\eta\nabla_{\lambda}L$. For the next input sample $X_{t}$, the updated SD-Calibrator is used for prediction. Using this single-sample single-step gradient descent strategy, we achieve lightning-fast parameter updates.
For clarity, we provide a Algorithm workflow 3 and Pytorch-Style Pseudocode 4 in Appendix C.2.
Complexity Analysis. The primary focus here is the time complexity. The Streaming Memory Queue itself has an $\mathcal{O}(1)$ time complexity for enqueue and dequeue operations. The Lightning Gradient Update is performed only once for each incoming test sample. Each update involves: 1). Forward propagation of the backbone and calibrator (dominated by the computational cost $\mathcal{O}(NTlogT)$ of rFFT and irFFT) 2). Backward propagation of the calibrator (dominated by parameter cost $\mathcal{O}(NG)$).
Theoretical Analysis. We also show that this single update step leads to a controlled adjustment, ensuring that the calibrator makes progress on the newest sample it’s trained on, without causing erratic behavior, under standard assumptions. (Please refer to Proposition 2 in Appendix B).
5 Experiments
In this section,we conduct extensive experiments to answer the following research questions (RQs):
  • RQ1: Can ST-TTC have a consistent improvement on various types of models and datasets? Can ST-TTC outperform previous learning methods that leverage test data? (Effectiveness)
  • RQ2: Can ST-TTC effective in various real-world scenarios, including few-shot learning, long-term forecasting, and large-scale forecasting? (Universality)
  • RQ3: Can ST-TTC further enhance the performance of existing learning paradigms that utilize training data, such as OOD Learning and continual learning? (Flexibility)
  • RQ4: How does ST-TTC work? Which components or strategies are crucial? Are these components or strategies sensitive to parameters or design? (Mechanism & Robustness)
  • RQ5: What is the time and parameter cost of ST-TTC during test-time computation, and how does it compare to other advanced methods? (Efficiency & Lightweight)
5.1 Experimental Setup
Datasets. We employ publicly available benchmark datasets widely used in the literature to cover typical spatio-temporal forecasting scenarios in the traffic domain (PEMS-03, PEMS-04, PEMS-07, PEMS-08 song2020spatial ), the meteorological domain (KnowAir wang2020pm2 ), and the energy domain (UrbanEV li2025urbanev ). In addition, we also leverage the traffic-speed benchmark METR-LA li2018diffusion , the large-scale spatio-temporal benchmark LargeST liu2023largest , and dynamic-stream benchmarks (Energy-Stream, Air-Stream, PEMS-Stream chen2025eac ) to assess our methods across varied settings and learning paradigms. Unless otherwise specified, all datasets are chronologically split into training, validation and test sets in a 6 : 2 : 2 ratio. For more detailed description of each dataset, please see the Appendix D.1.
Baseline. For the default evaluation, we cover various widely used spatio-temporal backbones, which can be divided into three categories: (1) Transformer-based: STAEformer liu2023spatio and STTN xu2020spatial ; (2) Graph-based: GWNet wu2019graph and STGCN yu2018spatio ; (3) MLP-based: STID shao2022spatial and ST-Norm deng2021st . For the baselines that leverage test information, we cover three types: (1) popular test-time adaptation methods in vision: TTT-MAE gandelsman2022test and TENT wangtent ; (2) Online time series forecasting methods: OnlineTCN zinkevich2003online , FSNet pham2023learning and OneNet wen2023onenet ; (3) Comparable online spatio-temporal forecasting methods: CompFormer zhang2023test and DOST wang2024distribution . For the baselines on large-scale benchmarks, we use the efficient PatchSTG fang2025efficient as the backbone. For the baseline of OOD learning scenarios, we use the advanced STONE wang2024stone as the default method. For the continual learning scenario, we use EAC chen2025eac and STKEC wang2023knowledge as the default methods. We follow the default parameter settings of the models for all scenarios according to the corresponding literature. For details of each method, see Appendix D.2.
Protocol. Following prior benchmarks shao2024exploring , we employ a 12-to-12 forecasting protocol—using the previous 12 time steps to predict the next 12 steps and their mean—evaluated with mean absolute error (MAE), root mean square error (RMSE), and mean absolute percentage error (MAPE). For simplicity, all experiments share the same hyperparameters of our ST-TTC : the calibration module learning rate $lr$ is set to 1e-4, the memory-queue sample count $n$ used for updating is 1, and the number of groups $m$ to 4. To ensure fairness, each experiment is repeated five times, with results reported as mean ± standard deviation (denoted in gray ±). More protocol details, see Appendix D.3.
5.2 Effectiveness Study (RQ1)
Consistent Effectiveness. Table LABEL:tab:rq1_consistent_performance_12 presents the results of our method for 12-step future prediction across six models on six public datasets. The  column denotes the results of standard testing, while the  column indicates results obtained with our proposed ST-TTC approach. The best results in the  and  columns are highlighted in bold blue and pink fonts, respectively. We also compute the relative improvement, denoted by the $\Delta$ column. Based on these results, we make the following observations: ❶ The application of our test-time computation method, ST-TTC , consistently yields performance gains across various backbone architectures and dataset combinations. ❷ From a model-centric perspective, our approach can further enhance the performance of even the top-performing methods across different metrics and datasets. ❸ From a data-centric perspective, UrbanEV shows more significant relative improvement, likely due to its more pronounced distribution shift.
Table 2: Performance comparison of different models w/ and w/o ST-TTC on common benchmarks.
Models Transformer-based Graph-based MLP-based
STAEformer liu2023spatio STTN xu2020spatial GWNet wu2019graph STGCN yu2018spatio STID shao2022spatial ST-Norm deng2021st
w/ ST-TTC $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%)
PEMS-03 MAE $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${16.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 1.47$ $18.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ ${17.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}}$ $\downarrow 1.32$ 16.73±0.41 16.42±0.19 $\downarrow 1.85$ $18.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}$ ${18.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}}$ $\downarrow 1.79$ $17.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${17.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}}$ $\downarrow 1.09$ $17.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ ${17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}}$ $\downarrow 1.39$
RMSE $29.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ ${29.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.58}}}$ $\downarrow 1.67$ $31.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.55}}$ ${30.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.37}}}$ $\downarrow 1.74$ 28.48±0.48 27.90±0.13 $\downarrow 2.04$ $31.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.94}}$ ${31.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.86}}}$ $\downarrow 2.02$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ ${28.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 1.24$ $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${28.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 1.95$
MAPE(%) 15.82±0.22 15.82±0.20 $\downarrow 0.00$ $18.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ ${18.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.06}}}$ $\downarrow 2.28$ $16.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.60}}$ ${16.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}}$ $\downarrow 1.26$ $18.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.78}}$ ${18.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}}$ $\downarrow 1.16$ $17.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ ${17.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}}$ $\downarrow 0.63$ $17.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ ${16.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}}$ $\downarrow 1.63$
PEMS-04 MAE 19.48±0.05 19.33±0.06 $\downarrow 0.77$ $20.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${20.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 0.73$ $20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.37}}$ ${20.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.39}}}$ $\downarrow 0.39$ $20.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ ${20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}}$ $\downarrow 0.63$ $19.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${19.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}}$ $\downarrow 0.55$ $20.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${20.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}}$ $\downarrow 0.69$
RMSE 32.58±0.39 32.31±0.34 $\downarrow 0.83$ $33.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${32.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.97$ $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}$ ${32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}}$ $\downarrow 0.46$ $33.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ ${32.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 0.94$ $32.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ ${32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}}$ $\downarrow 0.40$ $33.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ ${32.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}}$ $\downarrow 1.27$
MAPE(%) 12.42±0.01 12.37±0.09 $\downarrow 0.40$ $14.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ ${14.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}}$ $\downarrow 1.42$ $14.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.47}}$ ${14.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.32}}}$ $\downarrow 0.07$ $14.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${14.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}}$ $\downarrow 0.78$ $12.78{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${12.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.47$ $13.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ ${13.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}}$ $\downarrow 0.29$
PEMS-07 MAE 21.67±0.16 21.41±0.13 $\downarrow 1.20$ $23.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ ${23.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.75}}}$ $\downarrow 0.94$ $22.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ ${22.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}}$ $\downarrow 0.71$ $24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${23.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 1.77$ $21.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ ${21.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}}$ $\downarrow 0.78$ $22.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ ${22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 0.84$
RMSE $37.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ ${37.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.48}}}$ $\downarrow 1.20$ $37.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.91}}$ ${37.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}}$ $\downarrow 0.83$ $36.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${36.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}}$ $\downarrow 0.68$ $39.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ ${38.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 1.73$ 36.24±0.06 36.00±0.06 $\downarrow 0.66$ $38.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ ${37.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}}$ $\downarrow 0.97$
MAPE(%) 8.92±0.05 8.87±0.06 $\downarrow 0.56$ $10.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}$ ${9.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}}$ $\downarrow 0.99$ $9.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${9.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}}$ $\downarrow 0.41$ $10.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}$ ${10.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}}$ $\downarrow 1.80$ $9.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${9.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 0.55$ $9.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.56}}$ ${9.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.33}}}$ $\downarrow 0.80$
PEMS-08 MAE 14.84±0.09 14.73±0.08 $\downarrow 0.74$ $17.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.70$ $16.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ ${16.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 0.55$ $17.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ ${17.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 0.92$ $15.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${15.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.58$ $16.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${16.56{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}}$ $\downarrow 0.78$
RMSE 25.61±0.17 25.49±0.16 $\downarrow 0.47$ $27.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}$ ${26.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}}$ $\downarrow 0.52$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${26.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 0.34$ $27.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ ${27.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}}$ $\downarrow 0.65$ $25.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ ${25.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}}$ $\downarrow 0.39$ $26.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${26.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.52$
MAPE(%) 9.37±0.04 9.32±0.06 $\downarrow 0.53$ $11.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ ${11.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.28}}}$ $\downarrow 0.62$ $10.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.34}}$ ${10.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 1.46$ $11.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.36}}$ ${11.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}}$ $\downarrow 0.86$ $9.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ ${9.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.61$ $11.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.89}}$ ${11.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}}$ $\downarrow 2.36$
KnowAir MAE $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ ${17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 0.41$ $17.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.47$ 17.03±0.07 16.94±0.07 $\downarrow 0.53$ $17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${16.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.41$ $18.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ ${17.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.50$ $17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ ${17.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.35$
RMSE $26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${26.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}}$ $\downarrow 0.27$ $26.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.19$ 26.12±0.15 26.05±0.14 $\downarrow 0.27$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${26.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}}$ $\downarrow 0.27$ $27.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${27.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.22$ $26.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.23$
MAPE(%) $62.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.62}}$ ${61.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.53}}}$ $\downarrow 0.77$ $64.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.24}}$ ${63.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}}$ $\downarrow 1.51$ $64.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${63.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.28}}}$ $\downarrow 1.38$ $63.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ ${62.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.84}}}$ $\downarrow 0.67$ $70.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.44}}$ ${69.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.35}}}$ $\downarrow 1.33$ 60.76±0.34 60.50±0.63 $\downarrow 0.43$
UrbanEV MAE $2.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.70$ $3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${2.99{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 1.64$ $2.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 1.38$ $3.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${3.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 1.82$ 2.83±0.01 2.79±0.01 $\downarrow 1.41$ $3.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 1.62$
RMSE $5.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ ${4.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.40$ $5.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${5.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}}$ $\downarrow 1.18$ $4.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 1.23$ $5.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${5.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}}$ $\downarrow 1.95$ 4.74±0.02 4.67±0.02 $\downarrow 1.48$ $5.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 1.69$
MAPE(%) 27.14±0.46 26.75±0.58 $\downarrow 1.44$ $28.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ ${28.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}}$ $\downarrow 1.84$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.33}}$ ${28.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}}$ $\downarrow 2.16$ $31.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ ${31.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.01}}}$ $\downarrow 0.70$ $27.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ ${27.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.43}}}$ $\downarrow 1.63$ $29.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}$ ${29.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}}$ $\downarrow 0.91$
Table 3: Performance comparison of the advanced method with ST-TTC on METR-LA benchmark.
Method MAE RMSE w/o Training Set w/o Modifying Backbone
The training / validation / test set split used below is 70% / 10% / 20%.
TTT-MAE gandelsman2022test 3.47±0.03 7.43±0.05
TENT wangtent 4.84±0.08 8.53±0.10
CompFormer zhang2023test 3.46±0.02 7.19±0.08
ST-TTC 3.46±0.01 7.21±0.01
The training / validation / test set split used below is 20% / 5% / 75%.
OnlineTCN zinkevich2003online 4.78±0.03 8.70±0.04
FSNet pham2023learning 5.79±0.24 11.06±0.24
OneNet wen2023onenet 4.94±0.03 8.80±0.06
DOST wang2024distribution 4.38±0.02 8.26±0.03
ST-TTC 3.77±0.07 7.75±0.13
Competitive Effectiveness. We further compare our method against various advanced approaches that can leverage test-time information. Since the official source code of CompFormer and DOST is not available and uses additional data information, it leads to an unfair comparison. Nevertheless, we still include all reported METR-LA benchmark values (indicated with *) using a unified GWNet backbone, categorized into regular and online settings as presented in Table 3, based on their respective papers. Additionally, we implemented the popular TTT-MAE method as a surrogate for the unavailable TTT-ST method. Our observations are as follows: ❶ For the regular setting, our method achieves competitive results with more stable standard deviations. While other methods like CompFormer demonstrate similar performance, they often utilize more training information and computational resources. ❷ In the online setting, our method significantly outperforms existing approaches without requiring more complex model architecture modifications.
5.3 Universality Study (RQ2)
To demonstrate the universality of ST-TTC across diverse real-world scenarios, we explore various forecasting scenarios in the literature, including few-shot yuan2024spatio , long-term shao2022long , and large-scale han2024bigst .
Few-Shot Scenario. To simulate limited training data, we retrained models using only the first 10% of existing training sets to investigate a more common and challenging few-shot scenario. Figure 2 shows the relative performance gains with ST-TTC (For full results, please refer Table LABEL:tab:rq2_fewshot_12 in the Appendix). We observe: ❶ ST-TTC provides more significant improvements in the few-shot setting compared to the full-shot case in Table LABEL:tab:rq1_consistent_performance_12, with about half exceeding 2%. ❷ KnowAir shows the largest gain compared to other datasets, likely because its four-year long period leads to a substantial test distribution shift in the few-shot scenario, where our method adapts well.
Figure 2: Relative improvements of different models w/ ST-TTC in the few-shot setting.
Long-Term Scenario. In real-world scenarios, long-term forecasting helps to further plan future decisions. We predicted 24 future steps from 24 past steps to explore more complex temporal changes.
Figure 3: Left: relative improvement of long-term setting. Right: visualization study of PEMS-08.
As shown in Figure 3, we present the relative performance improvement of the advanced STID model with our ST-TTC method, and give a test set prediction visualization case on the PEMS-08 dataset (see Figure 9 in the Appendix for more examples). Our observations include: ❶ ST-TTC consistently improves long-term forecasting, even more than short-term (Table LABEL:tab:rq1_consistent_performance_12), likely due to more learnable information in longer windows. ❷ As the pink and orange box shows, our method learns test-time history, capturing both the global traffic decline and local fluctuations, leading to effective calibration.
Figure 4: Performance on LargeST.
Large-Scale Scenario. Beyond current regional datasets, state or national-level spatio-temporal forecasting can involve tens of thousands of stations and longer time frames. We explore large-scale scenarios using the popular LargeST benchmark (comprising SD, GBA, GLA, and CA subsets). Figure 4 illustrates the 12-step prediction performance gains of the state-of-the-art efficient spatio-temporal model PatchSTG fang2025efficient with our ST-TTC , along with a comparison of inference time complexity (For full results, see Table 7 in Appendix). We observe: ❶ Our ST-TTC consistently yields further performance improvements across all datasets, even surpassing the improvement of the second-best baseline over the PatchSTG on some datasets. ❷ The additional inference time is at most 3.82 minutes, which is a clear advantage for the achieved performance gains compared to the training time cost of up to 14 hours.
5.4 Flexibility Study (RQ3)
To illustrate the flexibility of ST-TTC in accommodating existing learning paradigms, we explore its integration with two training data-leveraging paradigms: OOD learning and Continual Learning.
Figure 5: Relative improvement using our ST-TTC method in the OOD learning setting.
OOD Learning Setting. Following prior work wang2024stone , we use the SD dataset to simulate spatio-temporal shift. For the temporal dimension, we use 1-8/2019, 9-10/2019, and 11-12/2020 for training, validation, and testing, respectively. For the spatial dimension, we randomly mask 10% of nodes in the test set and consider three proportions of new nodes (10% / 15% / 20%) relative to the training node to mimic varying degrees of shift. In Figure 5, we present the 12-step average prediction performance gains of the advanced OOD learning model STONE with our ST-TTC , evaluated on all nodes and new nodes to demonstrate generalizability and scalability (Full results in Table 8). We observe that: ❶ The STONE model with ST-TTC consistently achieves performance benefits, significantly outperforming all previous settings, indicating that existing OOD models are still insufficient for true OOD generalization, while our method is highly effective. ❷ For both all and new nodes, our improvements become more pronounced as the shift increases, further demonstrating our effectiveness in handling both generalizability and scalability in challenging scenarios.
Continual Learning Setting. Following prior work chen2025eac , we used multi-period streaming spatio-temporal data to examine our ST-TTC ’s integration with continual learning method. Table 4 shows the improved 12-step forecasting of advanced continual learning models EAC and STKEC with our ST-TTC . We observed: ❶ Consistent performance gains for both models across all datasets; STKEC with ST-TTC even achieved comparable performance to best model EAC. (2) Energy-Stream achieves significant improvement over other datasets, as the ST-TTC effectively learns and calibrates temporal changes, as shown by the frequency analysis (drastic shift changes) in Figure 6.
Table 4: Performance comparison in continual learning setting.
Methods w/ ST-TTC Air-Stream PEMS-Stream Energy-Stream
MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
EAC $24.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $38.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ $31.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $14.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ $20.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $5.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $50.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.60}}$
$23.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $37.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ $31.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $14.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $23.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ $20.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ $3.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ $3.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.00}}$ $39.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}$
$\Delta$downarrow 2.5% $\downarrow$ 1.9% $\downarrow$ 1.1% $\downarrow$ 1.4% $\downarrow$ 1.2% $\downarrow$ 1.4% $\downarrow$ 32.6% $\downarrow$ 27.8% $\downarrow$ 21.6%
STKEC $25.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.05}}$ $40.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $33.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.64}}$ $16.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $26.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $22.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $5.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $52.40{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$
$24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $39.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $31.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $16.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $21.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $4.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $43.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$
$\Delta$downarrow 4.6% $\downarrow$ 2.7% $\downarrow$ 4.7% $\downarrow$ 1.2% $\downarrow$ 1.3% $\downarrow$ 2.0% $\downarrow$ 29.2% $\downarrow$ 25.2% $\downarrow$ 17.5%
Figure 6: Energy-Stream’s Shift.
5.5 Mechanism & Robustness Study (RQ4)
Figure 7: Left: Strategy comparison. Middle: Effect of $lr$ w.r.t $m$ Right: Effect of $n$ w.r.t $s$.
We follow the OOD setup (challenging setting with 20% new nodes) to evaluate our ST-TTC .
Strategy Study. We compare different strategies: 1) simple nonlinear time domain calibration (Time), 2) learning only phase or amplitude modulation factors (Pha. / Amp.), 3) node-share modeling (Node), and (4) random selection or retrieval of the most similar samples (Rand. / Sim.). As shown in Figure 7 left, we observe: ❶ Frequency-domain calibration significantly outperforms time-domain calibration, with amplitude modulation being the primary contributor; ❷ Sharing nodes leads to performance degradation due to spatial heterogeneity in spatio-temporal data; ❸ Random sample selection reduces performance, and retrieving similar samples offers negligible gains while incurring higher computational cost. Our proposed update strategy is already near-optimal.
Parameter study. We analyze the sensitivity of two parameter groups. As shown in the middle and right of Figure 7: ❶ Higher learning rates and fewer groups generally lead to poorer performance, likely due to limited parameter capacity hindering stable learning; ❷ Increasing the number of samples or update steps has minimal impact on performance (fluctuations < 1%), but significantly increases time cost, highlighting the rationale of our flash update mechanism.
5.6 Efficiency & Lightweight Study (RQ5)
Figure 8: time and memory.
Result Analysis. We use GWNet as the backbone and compare ST-TTC with other test-time adaptation methods on METR-LA in terms of total inference time and memory usage. As shown in Figure 8, ST-TTC achieves the best overall efficiency (excluding the GWNet baseline), being 4.64× faster and reducing memory usage by 37.12% compared to the least efficient method. These improvements are significant as they are much smaller than the sliding size (5 min.), thereby meeting the real-time requirements.
6 Conclusion
In this paper, we investigate the objectives of test-time computation in spatio-temporal forecasting and explore effective approaches for its implementation. We propose ST-TTC , a noval paradigm that uses a flash gradient update with streaming memory queue to learning a spectral-domain calibrator via phase-amplitude modulation, effectively addressing non-stationary errors. Extensive experiments confirm its effectiveness, universality, and flexibility. In future work, we aim to explore how to enhance the internal computational capacity of spatio-temporal foundation models during test time.
Supplementary material
Learning with Calibration: Exploring
Test-Time Computing of Spatio-Temporal Forecasting
[…]
Appendix A More Related Work
A.1 Spectral Domain Learning.
Many recent forecasting models leverage spectral (Fourier or wavelet) representations to capture periodic or multiscale patterns in spatio-temporal data. For example, PastNet [79] integrates a Fourier-domain convolutional operator to embed physical inductive biases, achieving state-of-the-art results in weather and traffic prediction. FourierGNN [85] builds a learnable Fourier-graph operator that conducts graph convolutions in the frequency domain, reducing convolutional complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$. Wavelet-based methods like WDNO [25] perform diffusion modeling in the wavelet domain to capture abrupt spatio-temporal changes and multi-resolution features. In the pure time-series setting, approaches such as FITS [84] interpolate in the complex Fourier domain and discard negligible high-frequency components to maintain accuracy with very few parameters, and TimeKAN [26] explicitly decomposes multivariate series into multiple frequency bands using Kolmogorov–Arnold networks. These works demonstrate how frequency-domain learning can improve forecasting efficiency and accuracy by isolating dominant spectral components. Different from these methods, our method combines spectral domain feature extraction with calibration-aware test-time computation to achieve reliable and calibrated forecasts even under changing conditions.
A.2 Online Learning for Forecasting.
Traditional online forecasting methods include adaptive filters like Kalman filters and recursive least squares that update linear models on streaming data. Recently, deep-learning approaches have been proposed to handle nonstationarity in an online fashion. For instance, FSNet [55] implements a complementary “fast and slow” learning system: a fast-adapting component for sudden pattern changes and a slow memory component for repeating trends. OneNet [78] runs two parallel neural forecasters (one modeling temporal dependencies, one modeling cross-variable dependencies) and uses reinforcement learning to dynamically weight their predictions under concept drift. These methods continuously update model parameters or ensemble weights as new data arrive. A recent study [35] pointed out the information leakage problem of previous online time series prediction methods, where the model makes predictions and then evaluates them based on the historical time steps that have been back-propagated for parameter updates. By redefining the setting to focus on predicting unknown future steps and evaluating unobserved data points, they propose a two-stream framework for online prediction, DSOF, which is conceptually similar to previous methods, generating predictions in a coarse-to-fine manner through a teacher-student model. Compared with these methods, we focus on the more difficult spatio-temporal predictions while not requiring complex network architecture design. Instead, we propose a calibration-aware framework that focuses on adjusting predictions online instead of learning predictions.
A.3 Test-Time Adaptation.
Recent test-time adaptation techniques can be grouped by their adaptation strategy. Entropy minimization methods adjust a trained model to increase prediction confidence on unlabeled test data. For example, [71] propose TENT, which adapts model parameters by minimizing the entropy of its predictions on each test batch and updating batch-normalization layers online. Feature alignment methods recalibrate feature distributions using test inputs; for instance, adaptive batch-normalization techniques re-estimate BN statistics on the target data to align feature distributions without labels. Self-supervised adaptation uses auxiliary tasks on the test data to refine the model. Test-Time Training [64, 18, 24, 76] converts each test input into a self-supervised learning problem (e.g. predicting image rotations) and updates model parameters before making a prediction. Similarly, SHOT [43] freezes the source classifier and updates the feature extractor on unlabeled target data using pseudo-labeling and information maximization. Each of these paradigms improves generalization under distribution shift without access to target labels. In contrast, unlike these methods that exploit self-supervisory information, our spatio-temporal prediction setting can use labels from historical test information, enabling explicit optimization of the objective at test time, ensuring real-time adaptivity.
Appendix B Theoretical Analysis
B.1 Output Perturbation Bound
Theorem 1 (Output Perturbation Bound).
Let $Y\in\mathbb{C}^{B\times N\times M}$ be the original frequency-domain representation of the backbone’s prediction $y\in\mathbb{R}^{B\times N\times T}$, and $y^{\prime}\in\mathbb{R}^{B\times N\times T}$ be the calibrated output. Suppose the amplitude and phase modulation parameters satisfy $|\lambda^{\alpha}_{g}|\leq\epsilon_{\alpha}$ and $|\lambda^{\phi}_{g}|\leq\epsilon_{\phi}$ for all groups $g\in\{1,\dots,G\}$. Then, the $\ell_{2}$-norm of the calibration error satisfies:
$\|y^{\prime}-y\|_{2}\leq(\epsilon_{\alpha}+\epsilon_{\phi})\|Y\|_{2},$
where $\|Y\|_{2}$ is the $\ell_{2}$-norm of $Y$.
Proof.
Let $\Delta Y=Y^{\prime}-Y$ denote the frequency-domain perturbation. For each group $g$, the calibrated spectrum is $Y^{\prime}_{g}=A_{g}(1+\lambda^{\alpha}_{g})e^{j(P_{g}+\lambda^{\phi}_{g})}$. Expanding $Y^{\prime}_{g}$ around $\lambda^{\alpha}_{g}=0,\lambda^{\phi}_{g}=0$, we approximate:
$Y^{\prime}_{g}\approx Y_{g}\left(1+\lambda^{\alpha}_{g}+j\lambda^{\phi}_{g}\right),$
where higher-order terms (e.g., $\lambda^{\alpha}_{g}\lambda^{\phi}_{g}$) are neglected under small $\epsilon_{\alpha},\epsilon_{\phi}$. Thus, the perturbation is:
$\Delta Y_{g}\approx Y_{g}(\lambda^{\alpha}_{g}+j\lambda^{\phi}_{g}).$
The $\ell_{2}$-norm of $\Delta Y$ is bounded by:
$\|\Delta Y\|_{2}^{2}=\sum_{g=1}^{G}\sum_{f\in\text{Group }g}|\Delta Y_{g,f}|^{2}\leq\sum_{g=1}^{G}(\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2})\sum_{f\in\text{Group }g}|Y_{g,f}|^{2}=(\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2})\|Y\|_{2}^{2}.$
By Parseval’s theorem [53], $\|y^{\prime}-y\|_{2}=\|\Delta Y\|_{2}$, hence:
$\|y^{\prime}-y\|_{2}\leq\sqrt{\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2}}\|Y\|_{2}\leq(\epsilon_{\alpha}+\epsilon_{\phi})\|Y\|_{2}.$
Remark 1.
This theorem guarantees that the calibration-induced perturbation is linearly bounded by the modulation parameters $\epsilon_{\alpha},\epsilon_{\phi}$. By constraining these parameters (e.g., via regularization during test-time adaptation), SD-Calibrator ensures the calibrated output does not deviate excessively from the original prediction, thereby avoiding overfitting to transient noise. The group-wise parameterization further reduces the effective degrees of freedom (from $\mathcal{O}(NM)$ to $\mathcal{O}(NG)$), inherently limiting the risk of over-parameterization.
B.2 Controlled Descent on Streaming Memory Queues
Assumption 1 (Lipschitz Continuous Gradient of the Loss).
The loss function $L_{k}(\lambda)=\mathcal{L}(g_{\lambda}(f_{\theta}(X_{o}^{(k)})),Y_{o}^{(k)})$ is differentiable with respect to $\lambda$, and its gradient $\nabla_{\lambda}L_{k}(\lambda)$ is Lipschitz continuous with constant $L_{c}>0$. That is, for any $\lambda_{a},\lambda_{b}$:
$\|\nabla_{\lambda}L_{k}(\lambda_{a})-\nabla_{\lambda}L_{k}(\lambda_{b})\|_{2}\leq L_{c}\|\lambda_{a}-\lambda_{b}\|_{2}$
According to the descent Lemma [48], this implies:
$L_{k}(\lambda_{b})\leq L_{k}(\lambda_{a})+\langle\nabla_{\lambda}L_{k}(\lambda_{a}),\lambda_{b}-\lambda_{a}\rangle+\frac{L_{c}}{2}\|\lambda_{b}-\lambda_{a}\|_{2}^{2}$
Assumption 2 (Bounded Gradient).
The norm of the gradient of the loss function with respect to the calibrator parameters $\lambda$ is bounded for any sample $(X_{o}^{(k)},Y_{o}^{(k)})$ from the queue and any reasonable parameter set $\lambda_{k}$:
$\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}\leq G_{max}$
for some constant $G_{max}>0$.
This is a common assumption, especially if the output of the calibrator and the true labels are within a certain range, and the calibrator $g_{\lambda}$ is well-behaved.
Proposition 2 (Controlled Descent on Streaming Memory Queues).
Let the above assumptions hold. For the $k$-th update step using the dequeued sample pair $(X_{o}^{(k)},Y_{o}^{(k)})$, if the learning rate $\eta$ satisfies $0<\eta<\frac{2}{L_{c}}$, then the single gradient descent step on the SD-Calibrator parameters $\lambda$ ensures a decrease in the loss function for that specific sample:
$L_{k}(\lambda_{k+1})\leq L_{k}(\lambda_{k})-\eta\left(1-\frac{L_{c}\eta}{2}\right)\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}$
Furthermore, the change in the calibrator parameters is bounded:
$\|\lambda_{k+1}-\lambda_{k}\|_{2}\leq\eta G_{max}$
Proof.
Let $L_{k}(\lambda)=\mathcal{L}(g_{\lambda}(f_{\theta}(X_{o}^{(k)})),Y_{o}^{(k)})$ be the loss for the $k$-th dequeued sample. The parameter update rule is $\lambda_{k+1}=\lambda_{k}-\eta\nabla_{\lambda}L_{k}(\lambda_{k})$.
From Assumption 1, we have:
$L_{k}(\lambda_{k+1})\leq L_{k}(\lambda_{k})+\langle\nabla_{\lambda}L_{k}(\lambda_{k}),\lambda_{k+1}-\lambda_{k}\rangle+\frac{L_{c}}{2}\|\lambda_{k+1}-\lambda_{k}\|_{2}^{2}$
Substitute $\lambda_{k+1}-\lambda_{k}=-\eta\nabla_{\lambda}L_{k}(\lambda_{k})$:
$L_{k}(\lambda_{k+1})\leq L_{k}(\lambda_{k})+\langle\nabla_{\lambda}L_{k}(\lambda_{k}),-\eta\nabla_{\lambda}L_{k}(\lambda_{k})\rangle+\frac{L_{c}}{2}\|-\eta\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}$
$L_{k}(\lambda_{k+1})\leq L_{k}(\lambda_{k})-\eta\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}+\frac{L_{c}\eta^{2}}{2}\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}$
Factor out $\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}$:
$L_{k}(\lambda_{k+1})\leq L_{k}(\lambda_{k})-\eta\left(1-\frac{L_{c}\eta}{2}\right)\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}$
For the loss to decrease (or stay the same if gradient is zero), we require the term $\eta\left(1-\frac{L_{c}\eta}{2}\right)\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}\geq 0$. Since $\eta>0$ and $\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}^{2}\geq 0$, we need $\left(1-\frac{L_{c}\eta}{2}\right)>0$. This implies $1>\frac{L_{c}\eta}{2}$, so $\frac{2}{L_{c}}>\eta$. Thus, if $0<\eta<\frac{2}{L_{c}}$, the loss $L_{k}(\lambda_{k+1})$ on the sample $(X_{o}^{(k)},Y_{o}^{(k)})$ is strictly reduced if $\nabla_{\lambda}L_{k}(\lambda_{k})\neq 0$.
For the bound on parameter change:
$\|\lambda_{k+1}-\lambda_{k}\|_{2}=\|-\eta\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}=\eta\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}$
Using Assumption 2, $\|\nabla_{\lambda}L_{k}(\lambda_{k})\|_{2}\leq G_{max}$:
$\|\lambda_{k+1}-\lambda_{k}\|_{2}\leq\eta G_{max}$
This completes the proof. ∎
Remark 2.
The proposition demonstrates that each single-step update is not arbitrary but moves the SD-Calibrator’s parameters $\lambda$ in a direction that reduces the prediction error on the specific historical sample $(X_{o}^{(k)},Y_{o}^{(k)})$ used for the update, provided the learning rate $\eta$ is chosen appropriately (i.e., small enough, specifically $\eta<2/L_{c}$). The condition on $\eta$ ensures that the update step does not overshoot. The second part, $\|\lambda_{k+1}-\lambda_{k}\|_{2}\leq\eta G_{max}$, shows that the magnitude of change in the parameters $\lambda$ during each update is bounded. This is crucial for preventing the calibrator from experiencing excessively large or erratic parameter shifts from one step to the next, which could lead to instability or overfitting to noisy individual samples.
Appendix C Method Details
C.1 Spectral Domain Calibrator
Algorithm Workflow. We summarize the algorithm workflow of Section 4.1 in Algorithm 1.
Algorithm 1 Spectral Domain Calibrator
0:  Pre-trained backbone $f_{\theta}$, Test input $x$, Horizon length $T$, Number of nodes $N$, Groups $G$
0:  Calibrated output $\hat{y}^{\mathrm{cal}}$
1:  Get the backbone predictions: $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N\times T}$
2:  Compute $M\leftarrow\tfrac{T}{2}+1$vartriangleright I: Spatial-aware Decomposition
3:  Apply real-to-complex FFT along time dimension for each node: $Y_{f}\;\leftarrow\;\mathrm{rFFT}(\hat{y})\;\in\;\mathbb{C}^{N\times M}$
4:  Decompose: $A\leftarrow|Y_{f}|,\quad P\leftarrow\angle Y_{f}$vartriangleright II: Group-wise Modulation
5:  for $g=1,\dots,G$ do
6:     Get group index: $start\leftarrow(g-1)\lfloor M/G\rfloor+1$,  $end\leftarrow\begin{cases}M&g=G\\ g\lfloor M/G\rfloor&\text{otherwise}\end{cases}$
7:     Get learnable offsets: $\lambda^{\alpha}_{g}\in\mathbb{R}^{N\times 1},\;\lambda^{\phi}_{g}\in\mathbb{R}^{N\times 1}$
[…]
9:     Reconstruct slice: $Y^{\prime}_{f}[:,start:end]\leftarrow A^{\prime}_{g}\,\odot\,e^{j\,P^{\prime}_{g}}$
[…]
11:  Inverse FFT: $\hat{y}^{\mathrm{cal}}\leftarrow\mathrm{irFFT}\bigl{(}Y^{\prime}_{f}\bigr{)}\;\in\;\mathbb{R}^{N\times T}$
12:  return  $\hat{y}^{\mathrm{cal}}$
Algorithm Pseudo-code. We further present Algorithm 1 in the form of pytorch pseudo code in Algorithm 2 for easy understanding.
Algorithm 2 PyTorch-style pseudocode: SD-Calibrator Class
{minted}
python class SD_Calibrator(nn.Module): """ Spectral Domain Calibrator with Phase-Amplitude Modulation """
def __init__(self, num_nodes, freq_bins, groups=4): """ Args: num_nodes: number of spatial nodes (N) freq_bins: number of frequency bins (M = T // 2 + 1) groups: number of frequency groups (G) """ super().__init__() self.groups = groups self.group_size = freq_bins // groups
# Learnable offsets for amplitude and phase: (G, N, 1) self.lambda_amp = nn.Parameter( torch.zeros(groups, num_nodes, 1) ) self.lambda_phi = nn.Parameter( torch.zeros(groups, num_nodes, 1) )
def forward(self, y_pred): """ Args: y_pred: prediction from backbone, shape (B, 1, N, T) B defaults to 1, because only one sample can be tested Returns: calibrated prediction, shape (B, 1, N, T) """ B, _, N, T = y_pred.shape y = y_pred[:, 0] # (B, N, T)
Yf = torch.fft.rfft(y, dim=-1) # (B, N, M) A = torch.abs(Yf) P = torch.angle(Yf)
Yf_corr = torch.zeros_like(Yf) for g in range(self.groups): start = g * self.group_size if g == self.groups - 1 end = T // 2 + 1 else end = (g + 1) * self.group_size
lam_a = self.lambda_amp[g].unsqueeze(0) # (1, N, 1) lam_p = self.lambda_phi[g].unsqueeze(0)
A_g = A[:, :, start:end] * (1 + lam_a) P_g = P[:, :, start:end] + lam_p
Yf_corr[:, :, start:end] = A_g * torch.exp(1j * P_g)
y_time = torch.fft.irfft(Yf_corr, n=T, dim=-1) return y_time.unsqueeze(1) # (B, 1, N, T)
C.2 Lightning Gradient Update
Algorithm Pseudo-code. We summarize the algorithm workflow of Section 4.2 in Algorithm 3.
Algorithm 3 Flash Gradient Update Mechanism
0:  Test spatio-temporal sample stream $\{x_{t}\}_{t=1}^{B}$, Pre-trained backbone $f_{\theta}$, Streaming memory queue $\mathcal{Q}$, Queue size $T$ (equal to horizon length)
0:  Spectral domain calibrator $g_{\theta}$, Prediction collection of test samples $\{\hat{y}_{t}^{cal}\}_{t=1}^{B}$
1:  Initialize Calibrator module $g_{\theta}=(\lambda^{\alpha},\lambda^{\phi})$, empty queue $\mathcal{Q}$
2:  for each timestep $t=1,2,...$ do
3:     Receive $x_{t}$, compute default prediction $\hat{y}_{t}=f_{\theta}(x_{t})$vartriangleright I: Streaming Memory Queue
4:     Use Algorithm 1 to obtain the calibration results: $\hat{y}_{t}^{cal}=g_{\theta}(\hat{y}_{t}^{cal};\lambda)$
5:     Record ground truth: $y_{t}$ (collected by the value of $x_{t}$, available $T$ time steps in the future)
[…]
7:     if len($\mathcal{Q}$) $>T$ then
8:        $(x_{o},y_{o})=\mathcal{Q}$.dequeue()
9:        Use Algorithm 1 to obtain the calibration results: $\hat{y}_{o}^{cal}=f_{\theta}(x_{o})$
10:        Update: $\lambda\leftarrow\lambda-\eta\nabla_{\lambda}L(y_{o},\hat{y}_{o}^{cal})$
11:     end if
12:  end for
13:  return  $g_{\theta}$, $\{\hat{y}_{t}^{cal}\}_{t=1}^{B}$
Algorithm Workflow. We further present Algorithm 3 in the form of pytorch pseudo code in Algorithm 4 for easy understanding.
Algorithm 4 PyTorch-style pseudocode: Flash Gradient Update Function
{minted}
python def st_ttc_test(self, test_loader, node_num, T, groups): """ Flash Gradient Update with Streaming Memory Queue """ SDC = SD_Calibrator(node_num, T//2+1, groups).to(self.device) optimizer = torch.optim.Adam(SDC.parameters(), lr=1e-4) loss_fn = self._select_criterion() SMQ, preds = Queue(maxsize=T), []
for x, y in test_loader: x, y = x.to(self.device), y.to(self.device) with torch.no_grad(): y_pred = self.model(x) y_corr = SDC(y_pred)
# Use y_corr for inference y_corr = self.scaler.inverse_transform(y_corr) preds.append(y_corr.cpu().detach().numpy())
SMQ.put((x, y)) if SMQ.full(): x_old, y_old = SMQ.get() with torch.no_grad(): y_pred_old = self.model(x_old)
SDC.train() y_corr_old = SDC(y_pred_old) y_corr_old = self.scaler.inverse_transform(y_corr_old)
loss = loss_fn(y_corr_old, y_old) loss.backward() optimizer.step() optimizer.zero_grad() SDC.eval() return preds
Appendix D Experimental Details
D.1 Datasets Details
Our experiments are carried out on 14 real-world datasets from diffrent domain. The statistics of these spatio-temporal datasets are shown in Table 5.
Table 5: Summary of datasets used for our experiments. Degree: the average degree of each node. Meta: the number of metadata associated with each node. Data Points: multiplication of nodes and frames. M: million ($10^{6}$).
Source Dataset Nodes Time Range Frames Sampling Rate Data Points
 [63] PEMS03 358 09/01/2018 – 11/30/2018 26,208 5 minutes 9.38M
PEMS04 307 01/01/2018 – 02/28/2018 16,992 5 minutes 5.22M
PEMS07 883 05/01/2017 – 08/06/2017 28,224 5 minutes 24.92M
PEMS08 170 07/01/2016 – 08/31/2016 17,856 5 minutes 3.04M
 [37] UrbanEV 275 09/01/2022 – 02/28/2023 4344 1 hour 1.19M
 [77] Know-Air 184 01/01/2015 – 12/31/2018 11688 3 hours 2.15M
 [38] METR-LA 207 03/01/2012 – 06/27/2012 34,272 5 minutes 7.09M
LargeST [46] CA 8,600 01/01/2019 – 12/31/2019 35,040 15 minutes 30.13M
GLA 3,834 01/01/2019 – 12/31/2019 35,040 15 minutes 13.43M
GBA 2,352 01/01/2019 – 12/31/2019 35,040 15 minutes 8.87M
SD 716 01/01/2019 – 12/31/2020 70,080 15 minutes 5.02M
 [8] Air-Stream 1087 $\rightarrow$ 1154 $\rightarrow$ 1193 $\rightarrow$ 1202 01/01/2016 - 12/31/2019 34065 1 hour 15.79M
PEMS-Stream 655 $\rightarrow$ 715 $\rightarrow$ 786 $\rightarrow$ 822 $\rightarrow$ 834 $\rightarrow$ 850 $\rightarrow$ 871 07/10/2011 - 09/08/2017 61,992 5 minutes 34.30M
Energy-Stream 103 $\rightarrow$ 113 $\rightarrow$ 122 $\rightarrow$ 134 Unknown (245 days) 34,560 10 minutes 1.63M
We follow the conventional practice [38] to define the graph topology for all spatio-temporal datasets except Know-Air. Specifically, we construct the adjacency matrix $A$ for each dataset using a threshold Gaussian kernel, defined as follows:
$A_{[ij]}=\begin{cases}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)&\text{if}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)\geq r~{}\text{and}~{}i\neq j\\ 0&\text{otherwise}\end{cases}$
where $d_{ij}$ represents the distance between sensors $i$ and $j$, $\sigma$ is the standard deviation of all distances, and $r$ is the threshold. We follow the recommended parameter settings in all corresponding papers.
For the KnowAir dataset, we follow the original paper [77] and calculate the correlation between nodes to construct the adjacency matrix. Intuitively, most aerosol pollutants are distributed within a certain range above the ground. In addition, the mountains along the two cities will hinder the transmission of pollutants to the PM2.5 direction. Based on these intuitions, we constrain the weights in the adjacency matrix by the following formula:
$A_{[ij]}=H(d_{\theta}-d_{ij})\cdot H(m_{\theta}-m_{ij}),\quad\text{where}$
$d_{ij}=||\rho_{i}-\rho_{j}||,\quad m_{ij}=\sup_{\lambda\in(0,1)}\{h(\lambda\rho_{i}+(1-\lambda)\rho_{j})-\max\{h(\rho_{i}),h(\rho_{j})\}\},$
where $\rho_{i}$ is the location (latitude, longitude) of node $i$, $h(\rho)$ is the height of location $\rho$, and |||||$\cdot$||| | ⋅ | | is the L2-norm of the vector. $H(\cdot)$ is the Heaviside step function, where $H(x)=1$ if and only if $x>0$. $d_{\theta}$ and $m_{\theta}$ are the distance and altitude thresholds, respectively. Specifically, we also set the distance threshold $d_{\theta}=300$ km and the altitude threshold $m_{\theta}=1200$ meters.
D.2 Baseline Details
In our paper, we cover various spatio-temporal forecasting methods under various learning paradigms. The following is a classification and brief introduction of these advanced methods:
Classical Learning Methods for Spatio-Temporal Forecasting.
  • STAEformer [45]: STAEformer is a spatial-temporal adaptive embedding transformer that makes vanilla transformer state-of-the-art for spatio-temporal forecasting. It introduces a novel architecture to effectively capture the dynamic spatial and temporal dependencies in spatio-temporal data. https://github.com/XDZhelheim/STAEformer
  • STTN [83]: STTN is a spatial-temporal transformer network designed for traffic flow forecasting. It leverages dynamic directed spatial dependencies and long-range temporal dependencies to enhance the accuracy of long-term traffic predictions. https://github.com/xumingxingsjtu/STTN
  • GWNet [80]: GWNet is a graph wavenet model for deep spatial-temporal graph modeling. It effectively captures the complex spatial and temporal patterns in spatio-temporal data using a combination of graph convolutional networks and dilated causal convolutions. https://github.com/nnzhan/Graph-WaveNet
  • STGCN [87]: STGCN is a spatio-temporal graph convolutional network framework for traffic forecasting. It integrates graph convolutional networks with temporal convolutional networks to model the spatial and temporal dependencies in traffic data. https://github.com/hazdzz/stgcn
  • STID [60]: STID is a simple yet effective baseline for spatio-temporal forecasting. It addresses the indistinguishability of samples in spatial and temporal dimensions by attaching spatial and temporal identity information, achieving competitive performance with concise and efficient models. https://github.com/GestaltCogTeam/STID
  • ST-Norm [13]: ST-Norm is a method that applies spatial and temporal normalization for multi-variate time series forecasting. It enhances the performance of forecasting models by normalizing the spatial and temporal features of the data. https://github.com/JLDeng/ST-Norm
Efficient Learning Methods for Large-Scale Spatio-Temporal Forecasting.
  • PatchSTG [15]: PatchSTG is an attention-based dynamic spatial modeling method that uses irregular spatial patching for efficient large-scale spatio-temporal forecasting. It reduces computational complexity by segmenting large-scale inputs into balanced and non-overlapped patches, capturing local and global spatial dependencies effectively. https://github.com/lmissher/patchstg
OOD Learning Methods for Spatio-Temporal Forecasting.
  • STONE [66]: STONE is a state-of-the-art spatio-temporal OOD learning framework that effectively models spatial heterogeneity and generates temporal and spatial semantic graphs. It introduces a graph perturbation mechanism to enhance the model’s environmental modeling capability for better generalization. https://github.com/PoorOtterBob/STONE-KDD-2024
Continual Learning Methods for Spatio-Temporal Forecasting.
  • EAC [8]: EAC is a state-of-the-art method for exploring the rapid adaptation of models in the face of dynamic spatio-temporal graph changes during supervised finetuning. It follows the principles of expand and compress to address the challenges of retraining models over new data and catastrophic forgetting. https://github.com/Onedean/EAC
  • STKEC [68]: STKEC is a continual learning framework for traffic flow prediction on expanding traffic networks. It introduces a pattern bank to store representative network patterns and employs a pattern expansion mechanism to incorporate new patterns from evolving networks without requiring historical graph data. https://github.com/wangbinwu13116175205/STKEC
In addition to these advanced spatio-temporal forecasting models, we also cover various competitive baselines that learn with test information, mainly in the following three categories:
Popular test-time training methods
  • TTT-MAE [18]: TTT-MAE is a test-time training method that uses masked autoencoders to adjust the model during inference. It helps improve the performance of the model on unseen data by effectively utilizing test-time information. We adapt it to the backbone model of the spatiotemporal network, which is divided into a feature extractor and a prediction head as well as a self-supervisory head. https://github.com/Rima-ag/TTT-MAE
  • TENT [71]: TENT is a method for adjusting the model at test time by normalizing the activation function to reduce the offset between the training distribution and the test distribution. It enhances the generalization ability of the model without retraining on labeled test data. Although it is theoretically designed mainly for the cross entropy loss function, that is, classification tasks, we can still directly apply it to our prediction scenarios. https://github.com/DequanWang/tent
Classical online time series forecasting methods
  • OnlineTCN [101]: OnlineTCN is an online learning method based on a time convolutional network. It can adapt to new data sequentially and is very suitable for real-time prediction applications where data arrives continuously. https://github.com/locuslab/TCN
  • FSNet [55]: FSNet proposes a fast and slow learning network for online time series prediction that can handle both sudden changes and repeated patterns. In particular, FSNet improves on a slowly learning backbone by dynamically balancing fast adaptation to recent changes and retrieval of similar old knowledge. FSNet implements this mechanism through the interaction between two complementary components of the adapter to monitor each layer’s contribution to missing events, and an associative memory that supports remembering, updating, and recalling repeated events. https://github.com/salesforce/fsnet
  • OneNet [78]: OneNet dynamically updates and combines two models, one focusing on modeling dependencies across time dimensions and the other focusing on cross-variable dependencies. The approach integrates reinforcement learning-based methods into a traditional online convex programming framework, allowing the two models to be linearly combined with dynamically adjusted weights, thereby addressing the main drawback of classical online prediction methods that are slow to adapt to concept drift. https://github.com/yfzhang114/OneNet
Advanced spatio-temporal forecasting methods using test information.
  • CompFormer [97]: CompFormer proposes a test-time compensated representation learning framework, including a spatiotemporal decomposed database and a multi-head spatial transformer model. The former component explicitly separates all training data along the time dimension according to periodic features, while the latter component establishes connections between recent observations and historical sequences in the database through a spatial attention matrix. This enables it to transfer robust features to overcome abnormal events
  • DOST [70]: DOST proposes a novel online continuous learning framework tailored to the characteristics of spatiotemporal data. DOSTadopts an adaptive spatiotemporal network equipped with variable independent adapters to dynamically address the unique distribution changes of each urban location. In addition, to adapt to the gradual nature of these transformations, a wake-sleep learning strategy is used, which intermittently fine-tunes the adapters during the online stage to reduce computational overhead.
D.3 Protocol Details
Metrics Detail. We use different metrics such as MAE, RMSE, and MAPE. Formally, these metrics are formulated as following:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,~{}~{}~{}\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},~{}~{}~{}\text{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|$
where $n$ represents the indices of all observed samples, $y_{i}$ denotes the $i$-th actual sample, and $\hat{y_{i}}$ is the corresponding prediction.
Parameter Detail. For the hyper-parameter settings of all baseline methods, we follow the parameter settings recommended by the corresponding references. For our paper, except for the robustness study section, all other experimental hyper-parameters are set uniformly: the learning rate $lr$ is 1e-4, and the number of groups $m$ is set to 4. All experiments are conducted on a Linux server equipped with a 1 × AMD EPYC 7763 128-Core Processor CPU (256GB memory) and 4 × NVIDIA RTX A6000 (48GB memory) GPUs. To carry out benchmark testing experiments, all baselines are set to run for a duration of 100$\sim$150 epochs by default (depends on the corresponding paper), with specific timings contingent upon the method with early stop mechanism. The number of early stopping steps is set to 10.
Appendix E More Results
E.1 Complete Results Table
We provide complete information of the experimental tables in the main text as Table LABEL:tab:rq2_fewshot_1278
Table 6: Performance comparison of different models w/ and w/o ST-TTC in the few-shot scenario.
Models Transformer-based Graph-based MLP-based
STAEformer [45] STTN [83] GWNet [80] STGCN [87] STID [60] ST-Norm [13]
w/ ST-TTC
PEMS-03 MAE $23.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$ 22.96±0.93 $21.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.93}}$ 21.15±0.92 $21.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.98}}$ 21.43±0.92 $21.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 21.28±0.54 $21.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}$ 21.57±0.24 $21.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ 20.74±0.28
RMSE $37.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ 37.10±1.59 $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.59}}$ 33.87±1.53 $34.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.44}}$ 34.28±1.40 $34.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.92}}$ 34.08±0.89 $35.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ 35.18±0.52 $33.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.45}}$ 33.07±0.35
MAPE(%) $21.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.21}}$ 21.23±1.03 $21.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.35}}$ 21.13±1.28 $21.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.14}}$ 19.70±0.57 $22.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ 21.59±0.77 $21.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.86}}$ 21.27±0.73 $22.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.13}}$ 22.04±2.22
PEMS-04 MAE $35.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.25}}$ 34.57±4.08 $29.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.37}}$ 29.44±0.31 $33.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.86}}$ 32.87±1.95 $29.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ 29.66±0.83 $29.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 29.49±0.65 $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ 30.31±0.14
RMSE $50.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.86}}$ 50.23±4.59 $44.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.75}}$ 43.89±0.81 $50.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.56}}$ 49.73±2.73 $45.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.26}}$ 45.47±1.28 $44.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ 44.62±1.10 $45.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 45.33±0.49
MAPE(%) $23.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.28}}$ 23.40±3.26 $23.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.27}}$ 22.54±0.71 $22.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.52}}$ 22.43±3.17 $20.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ 20.72±1.19 $22.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.77}}$ 22.70±1.72 $21.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 21.72±0.60
PEMS-07 MAE $30.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.47}}$ 29.70±0.39 $31.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ 31.22±0.69 $33.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.65}}$ 32.82±0.63 $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.72}}$ 31.88±0.77 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.00}}$ 30.91±1.05 $31.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 30.50±0.05
RMSE $47.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ 46.89±0.76 $46.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$ 45.96±0.93 $49.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 49.36±0.57 $48.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ 47.61±0.05 $47.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ 46.71±0.97 $47.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.43}}$ 46.54±0.42
MAPE(%) $13.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 13.53±0.23 $14.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ 14.45±0.13 $15.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.70}}$ 14.74±0.58 $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}$ 16.28±0.99 $15.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.15}}$ 15.03±1.20 $14.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 14.18±0.46
PEMS-08 MAE $36.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.31}}$ 35.47±6.03 $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.42}}$ 23.81±0.41 $26.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.85}}$ 25.94±0.97 $25.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.25}}$ 25.31±0.21 $24.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 23.75±0.28 $24.34{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 24.06±0.07
RMSE $54.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±10.46}}$ 52.43±8.37 $36.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.45}}$ 36.61±0.50 $40.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.95}}$ 40.51±1.11 $38.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ 37.80±0.10 $37.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.77}}$ 37.36±0.77 $37.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ 37.20±0.20
MAPE(%) $27.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±9.55}}$ 26.19±8.29 $18.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ 17.65±0.54 $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$ 16.52±1.25 $20.16{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.91}}$ 17.84±0.88 $15.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ 14.43±0.20 $15.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 14.99±0.29
KnowAir MAE $18.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 18.16±0.35 $20.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ 19.85±0.23 $19.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.32}}$ 19.04±0.29 $20.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ 20.09±0.25 $22.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.20}}$ 21.72±0.94 $21.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.39}}$ 20.92±0.36
RMSE $27.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 26.97±0.08 $28.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ 28.51±0.46 $27.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ 27.58±0.15 $29.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ 28.65±0.21 $30.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.03}}$ 29.69±0.83 $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ 28.86±0.41
MAPE(%) $72.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.17}}$ 70.14±4.89 $85.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.82}}$ 81.21±1.63 $80.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.43}}$ 78.49±5.40 $84.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.40}}$ 82.13±2.18 $102.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.60}}$ 95.69±5.08 $95.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.26}}$ 91.11±1.80
UrbanEV MAE $3.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ 3.36±0.07 $4.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 4.05±0.06 $3.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 3.88±0.09 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 4.10±0.10 $3.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 3.24±0.05 $4.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 4.75±0.11
RMSE $6.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ 6.05±0.16 $6.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 6.71±0.05 $6.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ 6.59±0.11 $6.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ 6.61±0.17 $5.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 5.42±0.08 $8.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 8.36±0.27
MAPE(%) $31.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ 31.33±0.47 $38.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.43}}$ 37.32±1.33 $35.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.24}}$ 34.95±1.10 $40.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.48}}$ 39.57±1.52 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 30.75±0.50 $43.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.28}}$ 42.26±1.25
Table 7: PatchSTG with ST-TTC in LargeST Benchmark.
Datasets Methods Horizon 3 Horizon 6 Horizon 12 Average
MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
SD PatchSTG 14.53 24.34 9.22 16.86 28.63 11.11 20.66 36.27 14.72 16.90 29.27 11.23
w/ ST-TTC 14.44 24.07 8.70 16.66 28.18 10.93 20.37 35.68 14.27 16.72 28.83 11.11
$\Delta$downarrow 0.6% $\downarrow$ 1.1% $\downarrow$ 5.6% $\downarrow$ 1.2% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 3.1% $\downarrow$ 1.1% $\downarrow$ 1.5% $\downarrow$ 1.1%
GBA PatchSTG 16.81 28.71 12.25 19.68 33.09 14.51 23.49 39.23 18.93 19.50 33.16 14.64
w/ ST-TTC 16.65 28.31 12.12 19.30 32.40 14.35 22.96 38.33 18.30 19.16 32.49 14.48
$\Delta$downarrow 1.0% $\downarrow$ 1.4% $\downarrow$ 1.1% $\downarrow$ 1.9% $\downarrow$ 2.1% $\downarrow$ 1.1% $\downarrow$ 2.3% $\downarrow$ 2.3% $\downarrow$ 3.3% $\downarrow$ 1.7% $\downarrow$ 2.0% $\downarrow$ 1.1%
GLA PatchSTG 15.84 26.34 9.27 19.06 31.85 11.30 23.32 39.64 14.60 18.96 32.33 11.44
w/ ST-TTC 15.78 26.08 9.15 18.76 31.29 11.21 22.86 38.89 14.35 18.69 31.78 11.36
$\Delta$downarrow 0.4% $\downarrow$ 1.0% $\downarrow$ 1.3% $\downarrow$ 1.6% $\downarrow$ 1.8% $\downarrow$ 0.8% $\downarrow$ 2.0% $\downarrow$ 1.9% $\downarrow$ 1.7% $\downarrow$ 1.4% $\downarrow$ 1.7% $\downarrow$ 0.7%
CA PatchSTG 14.69 24.82 10.51 17.41 29.43 12.83 21.20 36.13 16.00 17.35 29.79 12.79
w/ ST-TTC 14.59 24.61 10.40 17.14 28.97 12.51 20.76 35.38 15.54 17.10 29.31 12.53
$\Delta$downarrow 0.7% $\downarrow$ 0.8% $\downarrow$ 1.0% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 2.5% $\downarrow$ 2.1% $\downarrow$ 2.1% $\downarrow$ 2.9% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 2.0%
Table 8: Performance of spatio-temporal shift dataset SD-ratio(%) on all nodes and unknown new nodes at different spatio-temporal shift levels.
Dataset Horizon Methods All Node New Node
MAE RMSE MAPE(%) MAE RMSE MAPE(%)
SD-10% 12 STONE $40.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.24}}$ $58.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.67}}$ $45.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±9.30}}$ $46.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.02}}$ $66.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.05}}$ $47.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±8.02}}$
w/ ST-TTC $39.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.82}}$ $57.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.67}}$ $38.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.19}}$ $44.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.69}}$ $65.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.95}}$ $41.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.87}}$
$\Delta$downarrow 3.3% $\downarrow$ 1.7% $\downarrow$ 17.0% $\downarrow$ 3.5% $\downarrow$ 2.9% $\downarrow$ 12.0%
Avg. STONE $30.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ $42.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.38}}$ $39.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.03}}$ $32.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.77}}$ $46.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.87}}$ $40.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}$
w/ ST-TTC $28.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.04}}$ $41.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.16}}$ $28.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.32}}$ $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.63}}$ $44.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.32}}$ $30.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.54}}$
$\Delta$downarrow 6.3% $\downarrow$ 3.1% $\downarrow$ 27.3% $\downarrow$ 6.5% $\downarrow$ 4.8% $\downarrow$ 23.2%
SD-15% 12 STONE $35.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.46}}$ $52.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}$ $34.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.73}}$ $43.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.85}}$ $66.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.28}}$ $30.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}$
w/ ST-TTC $34.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $52.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.54}}$ $29.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $41.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ $63.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $27.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$
$\Delta$downarrow 1.3% $\downarrow$ 1.2% $\downarrow$ 12.5% $\downarrow$ 3.9% $\downarrow$ 4.4% $\downarrow$ 9.1%
Avg. STONE $28.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $41.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ $36.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.91}}$ $33.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.69}}$ $49.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ $29.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.00}}$
w/ ST-TTC $26.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ $39.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $24.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.00}}$ $30.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $46.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ $22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}$
$\Delta$downarrow 6.5% $\downarrow$ 3.4% $\downarrow$ 30.7% $\downarrow$ 7.7% $\downarrow$ 6.0% $\downarrow$ 24.4%
SD-20% 12 STONE $36.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ $53.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $41.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.23}}$ $63.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.11}}$ $37.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.29}}$
w/ ST-TTC $35.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $52.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ $30.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $40.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ $60.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.07}}$ $33.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.73}}$
$\Delta$downarrow 2.9% $\downarrow$ 2.8% $\downarrow$ 10.6% $\downarrow$ 3.7% $\downarrow$ 3.8% $\downarrow$ 9.8%
Avg. STONE $28.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ $41.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $34.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.15}}$ $31.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.95}}$ $46.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.60}}$ $36.35{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.23}}$
w/ ST-TTC $26.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $39.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $24.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ $28.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $43.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.32}}$ $26.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.49}}$
$\Delta$downarrow 7.6% $\downarrow$ 4.8% $\downarrow$ 28.6% $\downarrow$ 8.0% $\downarrow$ 6.0% $\downarrow$ 26.8%
E.2 Visualization Case
We provide more visualization examples of test set predictions to illustrate the effectiveness of our calibration, as shown in Figure 9
Figure 9: Show case of improvement in spatio-temporal forecasting through our ST-TTC .
Appendix F More Discussion
F.1 Limitation
In this paper, we propose a novel paradigm for spatio-temporal forecasting: test-time computing. We systematically investigate the effective objectives for test-time computation and how to implement efficient computation. Based on detailed analysis, we posit that the key to the former lies in designing a lightweight, spatially-aware calibrator, while the key to the latter lies in designing an effective, non-information-leaking calibration parameter learning mechanism. Based on these insights, we present our corresponding solution, ST-TTC . We first propose a spectral domain calibrator, which appends a lightweight correction module operating in the frequency domain after the backbone network. This module calibrates errors by learning small, node-specific amplitude and phase correction factors. Furthermore, coupled with a flash gradient update mechanism featuring a streaming memory queue, it ensures universal, rapid, and resource-efficient test-time computing. Although there are still many potential areas for improvement, given the superiority and generality of our ST-TTC , we believe this provides a pathway for future exploration of larger-scale and more effective test-time computation. While we have taken a small step in this direction, several limitations warrant attention:
❶ Our current study does not involve testing on spatio-temporal foundation models. The fundamental reason behind this is our belief that true spatio-temporal foundation models do not yet exist. Although some preliminary exploratory work has been done [89, 90, 40, 39, 88], they are far from achieving true zero-shot generalization. However, considering their future inevitability, we believe that further improving the paradigm of test-time computation, especially how to activate and scale the internal capabilities of spatio-temporal foundation models during testing, goes beyond the design philosophy of our proposed learning with calibration. Nevertheless, our experiments still provide some preliminary guidance and insights.
❷ It is undoubtedly encouraging that our current calibration mechanism is more effective in large-scale and out-of-distribution scenarios. However, for commonly used small spatio-temporal benchmark datasets, the performance improvement is not yet significant. Therefore, how to effectively improve the performance of test-time computation on small-scale spatio-temporal datasets still requires exploration, and we reserve further improvement efforts for future research.
❸ We observed in our experiments that utilizing a larger amount of test information that is more similar to the current test sample does not significantly affect the results. This is certainly beneficial for real-time efficiency requirements. However, considering our current efficiency is already sufficiently good, further exploration is needed on how to potentially slow down the test-time computing process to make it more scalable and improve forecasting effectiveness.
F.2 Future Work
Building upon the research direction presented in this paper, we envision future work encompassing two main aspects:
❶ Exploring how to integrate retrieval-augmented techniques to filter more effective learning samples from arbitrary external scenarios, thereby combining them with our test-time computation framework to optimize performance on small-scale datasets.
❷ Investigating the construction of real spatio-temporal foundation models that encapsulate internal compressed knowledge, and exploring how to activate this internal capability during test time.
Appendix G Broader Impacts
This paper aims to promote the real-world usability of spatio-temporal forecasting models. We propose a novel paradigm, namely test-time computing of spatio-temporal forecasting. This paradigm shows significant generalization, universality across multiple scenarios, multiple tasks, multiple learning paradigms, and scalability to improve performance, providing valuable insights for future research and application value for practitioners. This paper focuses mainly on scientific research and has no obvious negative impact on society.

### v2

@@
 Extensive experiments on real-world datasets demonstrate the effectiveness, universality, flexibility and efficiency of our proposed method.
+[…]
+[…]
 1 Introduction
-Spatio-temporal forecasting (STF) aims to predict the future state of dynamic systems from historical spatio-temporal observations and underpins many real-world applications, such as traffic flow forecasting guo2021learning , air quality forecasting nguyen2023climax , and energy consumption forecasting wang2019review .
-Although spatio-temporal neural networks shao2024exploring ; jin2023spatio ; jin2024survey , which couple spatial neural operators with temporal neural operators, have achieved remarkable progress on these tasks, their deployment in practical environments remains fraught with challenges.
-These observations, typically collected by sensors, are frequently corrupted by noise, outliers (e.g., spikes or dropouts due to hardware failure) xu2024spatiotemporal , and more commonly, non-stationary distribution shifts arising from sensor aging and seasonal patterns wang2024evaluating .
-To enhance generalization and performance, prior work has focused primarily on out-of-distribution (OOD) learning for ST data during the training phase: designing architectures that resist perturbations ji2025seeing ; ma2024stop ; wang2024stone ; zhou2024coms2t ; zhou2023maintaining ; xia2023deciphering , augmenting training data with noise or adversarial examples bai2025alleviating ; zhang2023automated ; liu2022practical , and introducing specialized loss functions or regularizers li2024flashst ; zhang2023spatial to counteract distribution drift.
-However, these methods share fundamental limitations: they assume that the training data sufficiently captures all future target domain invariance—a premise that is rarely valid in real-world settings.
-Concurrently, an emerging paradigm of continual fine-tuning chen2025eac ; yi2024get ; wang2024towards ; lee2024continual ; wang2023pattern ; wang2023knowledge ; kieu2024team ; chen2021trafficstream has become popular in spatio-temporal learning by continuously tuning the model to adapt to dynamic changes.
+Spatio-temporal forecasting (STF) aims to predict the future state of dynamic systems from historical spatio-temporal observations and underpins many real-world applications, such as traffic flow forecasting (Guo et al., 2021), air quality forecasting (Nguyen et al., 2023), and energy consumption forecasting (Wang et al., 2019).
+Although spatio-temporal neural networks (Jin et al., 2023, 2024; Shao et al., 2024b), which couple spatial neural operators with temporal neural operators, have achieved remarkable progress on these tasks, their deployment in practical environments remains fraught with challenges.
+These observations, typically collected by sensors, are frequently corrupted by noise, outliers (e.g., spikes or dropouts due to hardware failure) (Xu et al., 2024a), and more commonly, non-stationary distribution shifts arising from sensor aging and seasonal patterns (Wang et al., 2024d).
+To enhance generalization and performance, prior work has focused primarily on out-of-distribution (OOD) learning for ST data during the training phase: designing architectures that resist perturbations (Ji et al., 2025; Ma et al., 2024; Wang et al., 2024a; Xia et al., 2023), augmenting training data with noise or adversarial examples (Bai et al., 2025; Liu et al., 2022; Zhang et al., 2023a), and introducing specialized loss functions or regularizers (Li et al., 2024b; Zhang et al., 2023b) to counteract distribution drift.
+However, these methods share fundamental limitations: they assume that the training data sufficiently captures all future target domain invariance, a premise that is rarely valid in real-world settings.
+Concurrently, an emerging paradigm of continual fine-tuning (Chen and Liang, 2025; Chen et al., 2021; Lee and Park, 2024; Wang et al., 2024b, 2023a, 2023b) has become popular in spatio-temporal learning by continuously tuning the model to adapt to dynamic changes.
 Though promising, it still divides the target domain into multiple periods of training and testing and relies on period-specific training data to optimize model, thereby failing in data-scarse scenarios.
@@
 Setting Example Works Data Perspective Learning Perspective
-Source Target Train-Time Test-Time
-OOD Learning STONE wang2024stone , CaST xia2023deciphering $\langle\mathbf{X}^{s},\mathbf{Y}^{s}\rangle$ $\min_{f_{\theta}}\;\max_{e^{*}\in\mathcal{E}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{s},\mathbf{Y}^{s}\mid e^{*})}\bigl{[}L(f_{\theta}(x),y)\bigr{]}$
-Continual Fine-Tuning EAC chen2025eac , TrafficStream chen2021trafficstream $\langle\mathbf{X}^{t},\mathbf{Y}^{t}\rangle$ $\min_{f_{{\theta}^{\tau}}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t},\mathbf{Y}^{t}\mid e^{\tau})}\bigl{[}L(f_{{\theta}^{\tau}}(x),y)\bigr{]}$
-Test-Time Training TTT-ST chen2024test $\mathbf{X}^{s}$ $\mathbf{X}^{t}$
+Source Target Train‐Time Test‐Time
+OOD Learning STONE (Wang et al., 2024a), CaST (Xia et al., 2023) $\langle\mathbf{X}^{s},\mathbf{Y}^{s}\rangle$ $\min_{f_{\theta}}\;\max_{e^{*}\in\mathcal{E}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{s},\mathbf{Y}^{s}\mid e^{*})}\bigl[L(f_{\theta}(x),y)\bigr]$
+Continual Fine‐Tuning EAC (Chen and Liang, 2025), TrafficStream (Chen et al., 2021) $\langle\mathbf{X}^{t},\mathbf{Y}^{t}\rangle$ $\min_{f_{{\theta}^{\tau}}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t},\mathbf{Y}^{t}\mid e^{\tau})}\bigl[L(f_{{\theta}^{\tau}}(x),y)\bigr]$
+Test‐Time Training TTT-ST (Chen et al., 2024) $\mathbf{X}^{s}$ $\mathbf{X}^{t}$
 $\min_{f_{\theta}}\,\mathbb{E}_{(\tilde{x},x)\sim P(\mathbf{X}^{s}\mid e)}\bigl[L(f_{\theta}(\tilde{x}),x)\bigr]$
 $\min_{f_{\theta}}\,\mathbb{E}_{(\tilde{x},x)\sim P(\mathbf{X}^{t}\mid e)}\bigl[L(f_{\theta}(\tilde{x}),x)\bigr]$
-Online Continual Learning DOST wang2024distribution $\mathbf{X}^{t}$ $\min_{f_{\theta(\delta)}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L(f_{\theta(\delta)}(x),y)\bigr{]}$
-Test-Time Computing ST-TTC (Our) $\mathbf{X}^{t}$ $\min_{g_{\theta}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L\bigl{(}g_{\theta}(f_{\theta}(x)),y\bigr{)}\bigr{]}$
+Online Continual Learning DOST (Wang et al., 2024c) $\mathbf{X}^{t}$ $\min_{f_{\theta(\delta)}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl[L(f_{\theta(\delta)}(x),y)\bigr]$
+Test‐Time Computing ST-TTC (Our) $\mathbf{X}^{t}$ $\min_{g_{\theta}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl[L\bigl(g_{\theta}(f_{\theta}(x)),y\bigr)\bigr]$
 Figure 1: Conceptual visualization comparison of different spatio-temporal learning paradigms under test environment.
@@
 Our (c) Test-Time Computing method only requires a lightweight calibrator $g_{\theta}$, which is a seamless and lightweight plug-and-play module.
-Recently, leveraging test-time information has attracted widespread attention for its ability to significantly improve language model performance on complex reasoning tasks snell2025scaling ; akyurek2024surprising .
-In computer vision, this concept has already been extensively developed: test-time training (TTT) was first introduced by sun2020test , which defines an auxiliary self-supervised task applied to both training and test samples to better balance bias and variance gandelsman2022test .
-A similar idea was adapted to spatio-temporal forecasting in TTT-ST chen2024test .
-Unlike language and vision settings—where obtaining ground-truth labels for test samples at inference time is nearly impossible—STF benefits from label autocorrelation cressie2011statistics : each observation strongly depends on its predecessor, and training instances are constructed from sliding windows, which provide access to historical samples and their true labels.
-Moreover, this property makes STF also require timeliness shao2024stemo , that is, the additional computing time during inference must be less than the window-stride interval.
-A recent STF method, DOST wang2024distribution , explores online continual learning, which initially explored this direction.
+Recently, leveraging test-time information has attracted widespread attention for its ability to significantly improve language model performance on complex reasoning tasks (Akyürek et al., 2024; Snell et al., 2025).
+In computer vision, this concept has already been extensively developed: test-time training (TTT) was first introduced by (Sun et al., 2020), which defines an auxiliary self-supervised task applied to both training and test samples to better balance bias and variance (Gandelsman et al., 2022).
+A similar idea was adapted to spatio-temporal forecasting in TTT-ST (Chen et al., 2024).
+Unlike language and vision settings—where obtaining ground-truth labels for test samples at inference time is nearly impossible—STF benefits from label autocorrelation (Cressie and Wikle, 2011): each observation strongly depends on its predecessor, and training instances are constructed from sliding windows, which provide access to historical samples and their true labels.
+Moreover, this property makes STF also require timeliness (Shao et al., 2024a), that is, the additional computing time during inference must be less than the window-stride interval.
+A recent STF method, DOST (Wang et al., 2024c), explores online continual learning, which initially explored this direction.
 It uses historical test sample labels to dynamically adapt the modified model architecture.
@@
 Spatio-temporal sequences can be regarded as spatially extended multivariate time series.
-Although one can trivially apply multivariate forecasting methods nie2023time ; zeng2023transformers ; biller2003modeling ; box1970distribution independently at each location, such decoupling of spatial and temporal dependencies invariably yields suboptimal results shao2024exploring .
-Classical spatio-temporal forecasting method instead relies on shallow models or spatio-temporal kernels, including feature-based methods ohashi2012wind ; zheng2015forecasting , state space models cliff1975space ; pfeifer1980starima ; bahadori2014fast , and Gaussian process models flaxman2015machine ; senanayake2016predicting .
-Unfortunately, the overall nonlinearity of these models is limited, and the high complexity of computation and storage further hinders the availability of massive training instances shi2018machine .
-In recent years, spatio-temporal neural networks kumar2024spatio ; jin2024survey ; jin2023spatio have been widely adopted to learn the complex dynamics of such systems.
-Early work concentrated on devising neural operators to extract spatial or temporal correlation yu2018spatio ; wu2019graph ; shao2022spatial ; liu2023spatio ; fang2025efficient and on designing fusion architectures to integrate them li2018diffusion ; guo2019attention ; oreshkin2021fc ; cini2023taming ; deng2024disentangling .
-More recent efforts have explored domain-invariant representation learning ma2024stop ; wang2024stone ; xia2023deciphering and continual model adaptation chen2025eac ; wang2023pattern ; chen2021trafficstream to better accommodate unseen environmental shifts.
+Although one can trivially apply multivariate forecasting methods (Biller and Nelson, 2003; Box and Pierce, 1970; Nie et al., 2023; Zeng et al., 2023) independently at each location, such decoupling of spatial and temporal dependencies invariably yields suboptimal results (Shao et al., 2024b).
+Classical spatio-temporal forecasting method instead relies on shallow models or spatio-temporal kernels, including feature-based methods (Ohashi and Torgo, 2012; Zheng et al., 2015), state space models (Bahadori et al., 2014; Cliff and Ord, 1975; Pfeifer and Deutsch, 1980), and Gaussian process models (Flaxman, 2015; Senanayake et al., 2016).
+Unfortunately, the overall nonlinearity of these models is limited, and the high complexity of computation and storage further hinders the availability of massive training instances (Shi and Yeung, 2018).
+In recent years, spatio-temporal neural networks (Jin et al., 2023, 2024; Kumar et al., 2024) have been widely adopted to learn the complex dynamics of such systems.
+Early work concentrated on devising neural operators to extract spatial or temporal correlation (Fang et al., 2025; Liu et al., 2023a; Shao et al., 2022b; Wu et al., 2019; Yu et al., 2018) and on designing fusion architectures to integrate them (Cini et al., 2023; Deng et al., 2024; Guo et al., 2019; Li et al., 2018; Oreshkin et al., 2021).
+More recent efforts have explored domain-invariant representation learning (Ma et al., 2024; Zhou et al., 2024, 2023) and continual model adaptation (Chen and Liang, 2025; Kieu et al., 2024; Chen et al., 2021; Yi et al., 2024) to better accommodate unseen environmental shifts.
 However, these methods still depend exclusively on offline training data and thus cannot deliver truly timely and effective adaptation in real settings.
 Test-Time Computing.
-Test-time computation is inspired by the human cognition kahneman2011thinking , in which additional computational effort is allocated during inference to improve task performance.
-This insight has recently driven considerable interest in the nature language process community, fueled by the success of reasoning-augmented language models (e.g., o1 jaech2024openai and r1 guo2025deepseek ) that activate and adapt internal computations at test time via supervised fine-tuning or reinforcement learning (RL) zhang2025and .
-While the generalization properties of RL-based adaptation remain debated yue2025does , the notion of supervised learning on unlabeled test data dates back to “transductive learning” Gammerman98lbyt in the 1990s and has demonstrated empirical benefits vapnik1999nature ; bottou1992local .
-In the computer vision domain, this idea was formalized as Test-Time Training sun2020test , which attaches an auxiliary self-supervised head to enable online adaptation to each test instance—a paradigm subsequently generalized as test-time adaptation wangtent ; jangtest ; wang2022continual ; liang2025comprehensive .
+Test-time computation is inspired by the human cognition (Kahneman, 2011), in which additional computational effort is allocated during inference to improve task performance.
+This insight has recently driven considerable interest in the nature language process community, fueled by the success of reasoning-augmented language models (e.g., o1 (Jaech et al., 2024) and r1 (Guo et al., 2025)) that activate and adapt internal computations at test time via supervised fine-tuning or reinforcement learning (RL) (Zhang et al., 2025).
+While the generalization properties of RL-based adaptation remain debated (Yue et al., 2025), the notion of supervised learning on unlabeled test data dates back to “transductive learning” (Gammerman et al., 1998) in the 1990s and has demonstrated empirical benefits (Bottou and Vapnik, 1992; Vapnik, 1999).
+In the computer vision domain, this idea was formalized as Test-Time Training (Sun et al., 2020), which attaches an auxiliary self-supervised head to enable online adaptation to each test instance—a paradigm subsequently generalized as test-time adaptation (Jang et al., 2023; Liang et al., 2025; Wang et al., 2021, 2022).
 However, spatio-temporal forecasting has seen limited exploration of such techniques.
-TTT-ST chen2024test applies TTT-style auxiliary objectives during training and continues to update at inference, and DOST wang2024distribution further incorporates dynamic learning mechanisms within modified model architectures for test-time updates.
-In addition, some methods zhang2023test ; guo2024online are conceptually close to ours, such as CompFormer zhang2023test , which proposes a test-time compensated representation learning framework, but still requires access to additional training data.
+TTT-ST (Chen et al., 2024) applies TTT-style auxiliary objectives during training and continues to update at inference, and DOST (Wang et al., 2024c) further incorporates dynamic learning mechanisms within modified model architectures for test-time updates.
+In addition, some methods (Guo et al., 2024; Zhang et al., 2023c) are conceptually close to ours, such as CompFormer (Zhang et al., 2023c), which proposes a test-time compensated representation learning framework, but still requires access to additional training data.
 Notably, we formalize the test-time computing of spatio-temporal forecasting, and propose a unified learning-with-calibration framework that is general, lightweight, efficient, and effective for STF at test-time.
@@
 Given historical observations $X^{h}\in\mathbb{R}^{N\times T^{h}\times C}$ (and an optional spatial correlation graph $\mathcal{G}$ representing the spatial relationships of $N$ locations), spatio-temporal forecasting aims to learn a mapping $f_{\theta}:(X^{h},\mathcal{G})\longmapsto X^{f}\in\mathbb{R}^{N\times T^{f}\times C}$, where $X^{f}$ is the signal for the next $T^{f}$ time steps.
-In practice, according to shao2024exploring ; li2018diffusion , the feature to be predicted is usually only the target variable.
+In practice, according to (Li et al., 2018; Shao et al., 2024b), the feature to be predicted is usually only the target variable.
 Scenario Definition.
@@
 In this section, we introduce these two key components, respectively, from the perspective of what is computed and how it is computed.
-4.1 What to Compute? Spectral Domain Calibrator with Phase-Amplitude Modulation
+4.1 What to Compute ?? Spectral Domain Calibrator with Phase-Amplitude Modulation
 Motivation.
 Spatio-temporal data, such as traffic flow and air quality, often exhibit periodic patterns (e.g., daily or weekly cycles).
-However, in real-world deployments, these patterns are not stationary; they are dynamically influenced by various internal and external factors wang2024robust .
+However, in real-world deployments, these patterns are not stationary; they are dynamically influenced by various internal and external factors (Wang et al., 2024e).
 Such influences lead to non-stationarities manifesting as fluctuations in amplitude (e.g., increased or decreased traffic peaks due to seasonal changes) or phase shifts (e.g., peak hours are advanced or delayed due to traffic congestion).
-Pre-training models typically fit fixed periodic patterns during training, which makes them vulnerable to performance degradation under such persistent dynamic changes during inference wang2024evaluating .
+Pre-training models typically fit fixed periodic patterns during training, which makes them vulnerable to performance degradation under such persistent dynamic changes during inference (Wang et al., 2024d).
 Therefore, we argue that the goal of test-time computation is: how to design an effective calibrator that can efficiently capture such gradual systematic bias from the pattern to correct the prediction errors caused by non-stationarity, while avoiding overfitting to random noise?
 Key Challenges.
-While correction in the time domain is possible zhang2023test ; guo2024online , it often requires extensive parameterization, leading to increased model complexity and limited ability to capture evolving periodic structures.
-Moreover, the coupled structural and branching modules wang2024distribution ; chen2024test are prone to overfitting the random noise in the spatio-temporal evolution.
+While correction in the time domain is possible (Guo et al., 2024; Zhang et al., 2023c), it often requires extensive parameterization, leading to increased model complexity and limited ability to capture evolving periodic structures.
+Moreover, the coupled structural and branching modules (Chen et al., 2024; Wang et al., 2024c) are prone to overfitting the random noise in the spatio-temporal evolution.
 To address this, we propose calibration in the spectral domain, where periodic variations are more transparently expressed as changes in the amplitude and phase of specific frequency components.
@@
   • Group-wise Modulation.
-To ensure lightweight and balanced spectrum expression, we divide the $M$ frequency bins into $G$ contiguous groups of size $\lfloor M/G\rfloor$, and learn per-group, per-node amplitude and phase offsets $\lambda^{\alpha}\in\mathbb{R}^{G\times N\times 1},\lambda^{\phi}\in\mathbb{R}^{G\times N\times 1}$.
+To ensure lightweight and balanced spectrum expression, we divide the $M$ frequency bins into $G$ contiguous groups of size $\lfloor M/G\rfloor$, and learn per‐group, per‐node amplitude and phase offsets $\lambda^{\alpha}\in\mathbb{R}^{G\times N\times 1},\lambda^{\phi}\in\mathbb{R}^{G\times N\times 1}$ (Note: Both $\lambda^{\alpha}$ and $\lambda^{\phi}$ are initialized to 0 to avoid incorrect calibration of predictions before learning).
 For each group $g\in\{1,\dots,G\}$, we apply $A_{g}^{{}^{\prime}}=A_{g}\odot(1+\lambda^{\alpha}_{g}\bigr),P_{g}^{{}^{\prime}}=P_{g}+\lambda^{\phi}_{g},$ and reconstruct the spectrum as $Y_{f}^{{}^{\prime}}=\bigcup_{g=1}^{G}A^{{}^{\prime}}_{g}\odot e^{(j\,P^{{}^{\prime}}_{g})}$.
   • Inverse Transform.
-Finally, the calibrated time-domain signal is obtained by Inverser rFFT $\hat{y}_{cal}=\mathrm{irFFT}\bigl{(}Y_{f})\;\in\;\mathbb{R}^{B\times N\times T},$ along the frequency dimension.
-For clarity, we provide a Algorithm workflow 1 and Pytorch-Style Pseudocode 2 in Appendix C.1.
+Finally, the calibrated time‐domain signal is obtained by Inverser rFFT $\hat{y}_{cal}=\mathrm{irFFT}\bigl(Y_{f})\;\in\;\mathbb{R}^{B\times N\times T},$ along the frequency dimension.
+For clarity, we provide a Algorithm workflow 1 and Pytorch-Style Pseudocode LABEL:pytorch_sdc in Appendix C.1.
 Complexity Analysis.
-The full-spectrum parameterization learns independent amplitude and phase offsets for each of the $M=T/2+1$ frequency bins and $N$ nodes, totaling $2NM$ parameters.
-In contrast, our $G$-group design learns only $2NG$ parameters.
+The full‐spectrum parameterization learns independent amplitude and phase offsets for each of the $M=T/2+1$ frequency bins and $N$ nodes, totaling $2NM$ parameters.
+In contrast, our GG‐group design learns only $2NG$ parameters.
 Since $G$ is a constant and $M$ grows linearly with $T$, $G\ll M$ is usually the case.
@@
 Theoretical Analysis.
-We also provide a theoretical bound on the output perturbation induced by the SD-Calibrator, ensuring controlled deviation from the original prediction to prevent overfitting (Please refer to Theorem 1 and the proof in Appendix B).
-4.2 How to Compute? Flash Gradient Update with Streaming Memory Queue
+We also provide a theoretical approximate bound on the output perturbation induced by the SD-Calibrator, ensuring controlled deviation from the original prediction to prevent overfitting (Please refer to Theorem 1 and the proof in Appendix B).
+4.2 How to Compute ?? Flash Gradient Update with Streaming Memory Queue
 Motivation.
@@
 Key Challenges.
-Although retrieving similar sequences from historical training databases can partially compensate for prediction errors zhang2023test , this assumption is unrealistic, as only test-time information is available in our scenario.
-Moreover, selectively storing historical test samples via memory bank primarily serves to mitigate catastrophic forgetting in the backbone model wang2024distribution , which misaligns with the learning objective of our SD-calibrator.
+Although retrieving similar sequences from historical training databases can partially compensate for prediction errors (Zhang et al., 2023c), this assumption is unrealistic, as only test-time information is available in our scenario.
+Moreover, selectively storing historical test samples via memory bank primarily serves to mitigate catastrophic forgetting in the backbone model (Wang et al., 2024c), which misaligns with the learning objective of our SD-calibrator.
 To address this, we propose freezing the backbone and updating only the calibrator using recent test samples for efficient test-time computing.
-However, this strategy introduces two critical challenges: ❶ recent studies lau2025fast have shown that real-time updates may cause information leakage; and ❷ excessive updates can lead to overfitting of the calibration parameters and increased computational burden.
+However, this strategy introduces two critical challenges: ❶ recent studies (Lau et al., 2025) have shown that real-time updates may cause information leakage; and ❷ excessive updates can lead to overfitting of the calibration parameters and increased computational burden.
 The core problem thus becomes how to design a efficient calibration parameter learning mechanism without information leakage.
@@
 Using this single-sample single-step gradient descent strategy, we achieve lightning-fast parameter updates.
-For clarity, we provide a Algorithm workflow 3 and Pytorch-Style Pseudocode 4 in Appendix C.2.
+For clarity, we provide a Algorithm workflow 2 and Pytorch-Style Pseudocode LABEL:pytorch_fgu in Appendix C.2.
 Complexity Analysis.
@@
 Datasets.
-We employ publicly available benchmark datasets widely used in the literature to cover typical spatio-temporal forecasting scenarios in the traffic domain (PEMS-03, PEMS-04, PEMS-07, PEMS-08 song2020spatial ), the meteorological domain (KnowAir wang2020pm2 ), and the energy domain (UrbanEV li2025urbanev ).
-In addition, we also leverage the traffic-speed benchmark METR-LA li2018diffusion , the large-scale spatio-temporal benchmark LargeST liu2023largest , and dynamic-stream benchmarks (Energy-Stream, Air-Stream, PEMS-Stream chen2025eac ) to assess our methods across varied settings and learning paradigms.
+We employ publicly available benchmark datasets widely used in the literature to cover typical spatio-temporal forecasting scenarios in the traffic domain (PEMS-03, PEMS-04, PEMS-07, PEMS-08 (Song et al., 2020)), the meteorological domain (KnowAir (Wang et al., 2020)), and the energy domain (UrbanEV (Li et al., 2025)).
+In addition, we also leverage the traffic-speed benchmark METR-LA (Li et al., 2018), the large-scale spatio-temporal benchmark LargeST (Liu et al., 2023b), and dynamic-stream benchmarks (Energy-Stream, Air-Stream, PEMS-Stream (Chen and Liang, 2025)) to assess our methods across varied settings and learning paradigms.
 Unless otherwise specified, all datasets are chronologically split into training, validation and test sets in a 6 : 2 : 2 ratio.
@@
 Baseline.
-For the default evaluation, we cover various widely used spatio-temporal backbones, which can be divided into three categories: (1) Transformer-based: STAEformer liu2023spatio and STTN xu2020spatial ; (2) Graph-based: GWNet wu2019graph and STGCN yu2018spatio ; (3) MLP-based: STID shao2022spatial and ST-Norm deng2021st .
-For the baselines that leverage test information, we cover three types: (1) popular test-time adaptation methods in vision: TTT-MAE gandelsman2022test and TENT wangtent ; (2) Online time series forecasting methods: OnlineTCN zinkevich2003online , FSNet pham2023learning and OneNet wen2023onenet ; (3) Comparable online spatio-temporal forecasting methods: CompFormer zhang2023test and DOST wang2024distribution .
-For the baselines on large-scale benchmarks, we use the efficient PatchSTG fang2025efficient as the backbone.
-For the baseline of OOD learning scenarios, we use the advanced STONE wang2024stone as the default method.
-For the continual learning scenario, we use EAC chen2025eac and STKEC wang2023knowledge as the default methods.
+For the default evaluation, we cover various widely used spatio-temporal backbones, which can be divided into three categories: (1) Transformer-based: STAEformer (Liu et al., 2023a) and STTN (Xu et al., 2020); (2) Graph-based: GWNet (Wu et al., 2019) and STGCN (Yu et al., 2018); (3) MLP-based: STID (Shao et al., 2022b) and ST-Norm (Deng et al., 2021).
+For the baselines that leverage test information, we cover three types: (1) popular test-time adaptation methods in vision: TTT-MAE (Gandelsman et al., 2022) and TENT (Wang et al., 2021); (2) Online time series forecasting methods: OnlineTCN (Zinkevich, 2003), FSNet (Pham et al., 2023) and OneNet (Wen et al., 2023); (3) Comparable online spatio-temporal forecasting methods: CompFormer (Zhang et al., 2023c) and DOST (Wang et al., 2024c).
+For the baselines on large-scale benchmarks, we use the efficient PatchSTG (Fang et al., 2025) as the backbone.
+For the baseline of OOD learning scenarios, we use the advanced STONE (Wang et al., 2024a) as the default method.
+For the continual learning scenario, we use EAC (Chen and Liang, 2025) and STKEC (Wang et al., 2023a) as the default methods.
 We follow the default parameter settings of the models for all scenarios according to the corresponding literature.
@@
 Protocol.
-Following prior benchmarks shao2024exploring , we employ a 12-to-12 forecasting protocol—using the previous 12 time steps to predict the next 12 steps and their mean—evaluated with mean absolute error (MAE), root mean square error (RMSE), and mean absolute percentage error (MAPE).
+Following prior benchmarks (Shao et al., 2024b), we employ a 12-to-12 forecasting protocol—using the previous 12 time steps to predict the next 12 steps and their mean—evaluated with mean absolute error (MAE), root mean square error (RMSE), and mean absolute percentage error (MAPE).
 For simplicity, all experiments share the same hyperparameters of our ST-TTC : the calibration module learning rate $lr$ is set to 1e-4, the memory-queue sample count $n$ used for updating is 1, and the number of groups $m$ to 4.
@@
 Models Transformer-based Graph-based MLP-based
-STAEformer liu2023spatio STTN xu2020spatial GWNet wu2019graph STGCN yu2018spatio STID shao2022spatial ST-Norm deng2021st
+STAEformer (Liu et al., 2023a) STTN (Xu et al., 2020) GWNet (Wu et al., 2019) STGCN (Yu et al., 2018) STID (Shao et al., 2022b) ST-Norm (Deng et al., 2021)
 w/ ST-TTC $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%) $\Delta$(%)
-PEMS-03 MAE $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${16.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 1.47$ $18.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ ${17.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}}$ $\downarrow 1.32$ 16.73±0.41 16.42±0.19 $\downarrow 1.85$ $18.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}$ ${18.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}}$ $\downarrow 1.79$ $17.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${17.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}}$ $\downarrow 1.09$ $17.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ ${17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}}$ $\downarrow 1.39$
-RMSE $29.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ ${29.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.58}}}$ $\downarrow 1.67$ $31.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.55}}$ ${30.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.37}}}$ $\downarrow 1.74$ 28.48±0.48 27.90±0.13 $\downarrow 2.04$ $31.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.94}}$ ${31.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.86}}}$ $\downarrow 2.02$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ ${28.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 1.24$ $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${28.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 1.95$
-MAPE(%) 15.82±0.22 15.82±0.20 $\downarrow 0.00$ $18.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ ${18.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.06}}}$ $\downarrow 2.28$ $16.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.60}}$ ${16.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}}$ $\downarrow 1.26$ $18.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.78}}$ ${18.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}}$ $\downarrow 1.16$ $17.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ ${17.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}}$ $\downarrow 0.63$ $17.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ ${16.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}}$ $\downarrow 1.63$
-PEMS-04 MAE 19.48±0.05 19.33±0.06 $\downarrow 0.77$ $20.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${20.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 0.73$ $20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.37}}$ ${20.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.39}}}$ $\downarrow 0.39$ $20.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ ${20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}}$ $\downarrow 0.63$ $19.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${19.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}}$ $\downarrow 0.55$ $20.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${20.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}}$ $\downarrow 0.69$
-RMSE 32.58±0.39 32.31±0.34 $\downarrow 0.83$ $33.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${32.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.97$ $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.35}}$ ${32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}}$ $\downarrow 0.46$ $33.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ ${32.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 0.94$ $32.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ ${32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}}$ $\downarrow 0.40$ $33.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ ${32.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}}$ $\downarrow 1.27$
-MAPE(%) 12.42±0.01 12.37±0.09 $\downarrow 0.40$ $14.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ ${14.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}}$ $\downarrow 1.42$ $14.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.47}}$ ${14.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.32}}}$ $\downarrow 0.07$ $14.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${14.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}}$ $\downarrow 0.78$ $12.78{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${12.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.47$ $13.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ ${13.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}}$ $\downarrow 0.29$
-PEMS-07 MAE 21.67±0.16 21.41±0.13 $\downarrow 1.20$ $23.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ ${23.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.75}}}$ $\downarrow 0.94$ $22.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ ${22.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}}$ $\downarrow 0.71$ $24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${23.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 1.77$ $21.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ ${21.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}}$ $\downarrow 0.78$ $22.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ ${22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}}$ $\downarrow 0.84$
-RMSE $37.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ ${37.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.48}}}$ $\downarrow 1.20$ $37.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.91}}$ ${37.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}}$ $\downarrow 0.83$ $36.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${36.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}}$ $\downarrow 0.68$ $39.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ ${38.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 1.73$ 36.24±0.06 36.00±0.06 $\downarrow 0.66$ $38.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ ${37.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}}$ $\downarrow 0.97$
-MAPE(%) 8.92±0.05 8.87±0.06 $\downarrow 0.56$ $10.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}$ ${9.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}}$ $\downarrow 0.99$ $9.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${9.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}}$ $\downarrow 0.41$ $10.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}$ ${10.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}}$ $\downarrow 1.80$ $9.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${9.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 0.55$ $9.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.56}}$ ${9.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.33}}}$ $\downarrow 0.80$
-PEMS-08 MAE 14.84±0.09 14.73±0.08 $\downarrow 0.74$ $17.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.70$ $16.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ ${16.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}}$ $\downarrow 0.55$ $17.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ ${17.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 0.92$ $15.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${15.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.58$ $16.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${16.56{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}}$ $\downarrow 0.78$
-RMSE 25.61±0.17 25.49±0.16 $\downarrow 0.47$ $27.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}$ ${26.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}}$ $\downarrow 0.52$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${26.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}}$ $\downarrow 0.34$ $27.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ ${27.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}}$ $\downarrow 0.65$ $25.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ ${25.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}}$ $\downarrow 0.39$ $26.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${26.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.52$
-MAPE(%) 9.37±0.04 9.32±0.06 $\downarrow 0.53$ $11.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ ${11.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.28}}}$ $\downarrow 0.62$ $10.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.34}}$ ${10.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 1.46$ $11.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.36}}$ ${11.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}}$ $\downarrow 0.86$ $9.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ ${9.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.61$ $11.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.89}}$ ${11.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}}$ $\downarrow 2.36$
-KnowAir MAE $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ ${17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.18}}}$ $\downarrow 0.41$ $17.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.47$ 17.03±0.07 16.94±0.07 $\downarrow 0.53$ $17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${16.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.41$ $18.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ ${17.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 0.50$ $17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ ${17.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.35$
-RMSE $26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${26.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}}$ $\downarrow 0.27$ $26.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ ${26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}}$ $\downarrow 0.19$ 26.12±0.15 26.05±0.14 $\downarrow 0.27$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ ${26.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}}$ $\downarrow 0.27$ $27.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${27.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.22$ $26.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 0.23$
-MAPE(%) $62.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.62}}$ ${61.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.53}}}$ $\downarrow 0.77$ $64.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.24}}$ ${63.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}}$ $\downarrow 1.51$ $64.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ ${63.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.28}}}$ $\downarrow 1.38$ $63.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ ${62.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.84}}}$ $\downarrow 0.67$ $70.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.44}}$ ${69.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.35}}}$ $\downarrow 1.33$ 60.76±0.34 60.50±0.63 $\downarrow 0.43$
-UrbanEV MAE $2.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.70$ $3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ ${2.99{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 1.64$ $2.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}}$ $\downarrow 1.38$ $3.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ ${3.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}}$ $\downarrow 1.82$ 2.83±0.01 2.79±0.01 $\downarrow 1.41$ $3.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ ${3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 1.62$
-RMSE $5.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ ${4.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.40$ $5.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${5.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}}$ $\downarrow 1.18$ $4.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 1.23$ $5.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${5.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}}$ $\downarrow 1.95$ 4.74±0.02 4.67±0.02 $\downarrow 1.48$ $5.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 1.69$
-MAPE(%) 27.14±0.46 26.75±0.58 $\downarrow 1.44$ $28.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ ${28.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}}$ $\downarrow 1.84$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.33}}$ ${28.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}}$ $\downarrow 2.16$ $31.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ ${31.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.01}}}$ $\downarrow 0.70$ $27.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ ${27.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.43}}}$ $\downarrow 1.63$ $29.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}$ ${29.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}}$ $\downarrow 0.91$
-Table 3: Performance comparison of the advanced method with ST-TTC on METR-LA benchmark.
-Method MAE RMSE w/o Training Set w/o Modifying Backbone
-The training / validation / test set split used below is 70% / 10% / 20%.
-TTT-MAE gandelsman2022test 3.47±0.03 7.43±0.05
-TENT wangtent 4.84±0.08 8.53±0.10
-CompFormer zhang2023test 3.46±0.02 7.19±0.08
-ST-TTC 3.46±0.01 7.21±0.01
-The training / validation / test set split used below is 20% / 5% / 75%.
-OnlineTCN zinkevich2003online 4.78±0.03 8.70±0.04
-FSNet pham2023learning 5.79±0.24 11.06±0.24
-OneNet wen2023onenet 4.94±0.03 8.80±0.06
-DOST wang2024distribution 4.38±0.02 8.26±0.03
-ST-TTC 3.77±0.07 7.75±0.13
+PEMS-03 MAE $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $16.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$} $\downarrow 1.47$ $18.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.53}}$ $17.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.50}}$} $\downarrow 1.32$ 16.73±0.41 16.42±0.19 $\downarrow 1.85$ $18.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.35}}$ $18.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.35}}$} $\downarrow 1.79$ $17.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ $17.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$} $\downarrow 1.09$ $17.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.13}}$ $17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$} $\downarrow 1.39$
+RMSE $29.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.59}}$ $29.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.58}}$} $\downarrow 1.67$ $31.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.55}}$ $30.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.37}}$} $\downarrow 1.74$ 28.48±0.48 27.90±0.13 $\downarrow 2.04$ $31.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.94}}$ $31.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.86}}$} $\downarrow 2.02$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$ $28.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$} $\downarrow 1.24$ $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$ $28.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$} $\downarrow 1.95$
+MAPE(%) 15.82±0.22 15.82±0.20 $\downarrow 0.00$ $18.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.19}}$ $18.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.06}}$} $\downarrow 2.28$ $16.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.60}}$ $16.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.30}}$} $\downarrow 1.26$ $18.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.78}}$ $18.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$} $\downarrow 1.16$ $17.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$ $17.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$} $\downarrow 0.63$ $17.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.82}}$ $16.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.52}}$} $\downarrow 1.63$
+PEMS-04 MAE 19.48±0.05 19.33±0.06 $\downarrow 0.77$ $20.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $20.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$} $\downarrow 0.73$ $20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.37}}$ $20.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.39}}$} $\downarrow 0.39$ $20.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ $20.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$} $\downarrow 0.63$ $19.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $19.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$} $\downarrow 0.55$ $20.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $20.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$} $\downarrow 0.69$
+RMSE 32.58±0.39 32.31±0.34 $\downarrow 0.83$ $33.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $32.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$} $\downarrow 0.97$ $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.35}}$ $32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.38}}$} $\downarrow 0.46$ $33.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$ $32.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$} $\downarrow 0.94$ $32.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $32.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$} $\downarrow 0.40$ $33.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ $32.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}}$} $\downarrow 1.27$
+MAPE(%) 12.42±0.01 12.37±0.09 $\downarrow 0.40$ $14.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.76}}$ $14.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.41}}$} $\downarrow 1.42$ $14.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.47}}$ $14.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.32}}$} $\downarrow 0.07$ $14.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $14.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$} $\downarrow 0.78$ $12.78{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$ $12.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 0.47$ $13.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.13}}$ $13.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.24}}$} $\downarrow 0.29$
+PEMS-07 MAE 21.67±0.16 21.41±0.13 $\downarrow 1.20$ $23.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ $23.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.75}}$} $\downarrow 0.94$ $2.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}} 22$$2.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$} $\downarrow 0.71$ $24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}}$ $23.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.18}}$} $\downarrow 1.77$ $21.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $21.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$} $\downarrow 0.78$ $2.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}} 22$$2.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$} $\downarrow 0.84$
+RMSE $37.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.52}}$ $37.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.48}}$} $\downarrow 1.20$ $37.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.91}}$ $37.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$} $\downarrow 0.83$ $36.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}}$ $36.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$} $\downarrow 0.68$ $39.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$ $38.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$} $\downarrow 1.73$ 36.24±0.06 36.00±0.06 $\downarrow 0.66$ $38.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$ $37.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.41}}$} $\downarrow 0.97$
+MAPE(%) 8.92±0.05 8.87±0.06 $\downarrow 0.56$ $10.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.24}}$ $9.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.24}}$} $\downarrow 0.99$ $9.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $9.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$} $\downarrow 0.41$ $10.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.18}}$ $10.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$} $\downarrow 1.80$ $9.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $9.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$} $\downarrow 0.55$ $9.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.56}}$ $9.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.33}}$} $\downarrow 0.80$
+PEMS-08 MAE 14.84±0.09 14.73±0.08 $\downarrow 0.74$ $17.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$ $17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$} $\downarrow 0.70$ $16.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$ $16.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$} $\downarrow 0.55$ $33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.19}} 17.17$$17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$} $\downarrow 0.92$ $15.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ $15.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.58$ $16.69{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $16.56{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.04}}$} $\downarrow 0.78$
+RMSE 25.61±0.17 25.49±0.16 $\downarrow 0.47$ $27.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.30}}$ $26.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$} $\downarrow 0.52$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}}$ $26.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$} $\downarrow 0.34$ $27.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$ $27.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$} $\downarrow 0.65$ $25.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $25.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$} $\downarrow 0.39$ $26.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $26.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$} $\downarrow 0.52$
+MAPE(%) 9.37±0.04 9.32±0.06 $\downarrow 0.53$ $1.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}} 11$$1.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.28}}$} $\downarrow 0.62$ $10.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.34}}$ $10.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.18}}$} $\downarrow 1.46$ $1.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.36}} 11$$1.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.30}}$} $\downarrow 0.86$ $9.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$ $9.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 0.61$ $1.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.89}} 11$$1.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.38}}$} $\downarrow 2.36$
+KnowAir MAE $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.19}}$ $17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.18}}$} $\downarrow 0.41$ $17.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $17.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$} $\downarrow 0.47$ 17.03±0.07 16.94±0.07 $\downarrow 0.53$ $17.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $16.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 0.41$ $18.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ $17.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$} $\downarrow 0.50$ $17.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$ $17.01{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.35$
+RMSE $26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$ $26.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.19}}$} $\downarrow 0.27$ $26.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $26.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$} $\downarrow 0.19$ 26.12±0.15 26.05±0.14 $\downarrow 0.27$ $26.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$ $26.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$} $\downarrow 0.27$ $27.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ $27.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.22$ $26.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 0.23$
+MAPE(%) $62.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.62}}$ $61.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.53}}$} $\downarrow 0.77$ $64.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.24}}$ $63.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.40}}$} $\downarrow 1.51$ $64.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$ $63.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.28}}$} $\downarrow 1.38$ $63.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.97}}$ $62.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.84}}$} $\downarrow 0.67$ $70.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.44}}$ $69.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.35}}$} $\downarrow 1.33$ 60.76±0.34 60.50±0.63 $\downarrow 0.43$
+UrbanEV MAE $2.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ $2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.70$ $3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $2.99{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 1.64$ $2.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $2.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$} $\downarrow 1.38$ $3.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $3.23{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$} $\downarrow 1.82$ 2.83±0.01 2.79±0.01 $\downarrow 1.41$ $3.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ $3.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 1.62$
+RMSE $5.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$ $4.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.40$ $5.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $5.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$} $\downarrow 1.18$ $4.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 1.23$ $63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}} 5.5$$52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$} $\downarrow 1.95$ 4.74±0.02 4.67±0.02 $\downarrow 1.48$ $5.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 1.69$
+MAPE(%) 27.14±0.46 26.75±0.58 $\downarrow 1.44$ $28.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $28.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$} $\downarrow 1.84$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.33}}$ $28.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$} $\downarrow 2.16$ $31.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.97}}$ $31.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.01}}$} $\downarrow 0.70$ $27.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.62}}$ $27.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.43}}$} $\downarrow 1.63$ $29.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.57}}$ $29.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.57}}$} $\downarrow 0.91$
 Competitive Effectiveness.
@@
 Since the official source code of CompFormer and DOST is not available and uses additional data information, it leads to an unfair comparison.
-Nevertheless, we still include all reported METR-LA benchmark values (indicated with *) using a unified GWNet backbone, categorized into regular and online settings as presented in Table 3, based on their respective papers.
+Nevertheless, we still include all reported METR-LA benchmark values (indicated with *) using a unified GWNet backbone, categorized
+Table 3: Performance comparison of the advanced method w/ ST-TTC on METR-LA benchmark. means the results are statistically significant (t-test with p-value <0.01<0.01).
+Method MAE RMSE
+w/o Training
+Set
+w/o Modifying
+Backbone
+The training / validation / test set split used below is 70% / 10% / 20%.
+TTT-MAE (Gandelsman et al., 2022) 3.47±0.03 7.43±0.05
+TENT (Wang et al., 2021) 4.84±0.08 8.53±0.10
+CompFormer (Zhang et al., 2023c) 3.46±0.02 7.19±0.08
+ST-TTC 3.46±0.01 7.21±0.01
+The training / validation / test set split used below is 20% / 5% / 75%.
+OnlineTCN (Zinkevich, 2003) 4.78±0.03 8.70±0.04
+FSNet (Pham et al., 2023) 5.79±0.24 11.06±0.24
+OneNet (Wen et al., 2023) 4.94±0.03 8.80±0.06
+DOST (Wang et al., 2024c) 4.38±0.02 8.26±0.03
+ST-TTC 3.77±0.07 7.75±0.13
+into regular and online settings as presented in Table 3, based on their respective papers.
 Additionally, we implemented the popular TTT-MAE method as a surrogate for the unavailable TTT-ST method.
@@
 5.3 Universality Study (RQ2)
-To demonstrate the universality of ST-TTC across diverse real-world scenarios, we explore various forecasting scenarios in the literature, including few-shot yuan2024spatio , long-term shao2022long , and large-scale han2024bigst .
+To demonstrate the universality of ST-TTC across diverse real-world scenarios, we explore various forecasting scenarios in the literature, including few-shot (Yuan et al., 2024c), long-term (Shao et al., 2022a), and large-scale (Han et al., 2024).
 Few-Shot Scenario.
@@
 Figure 3: Left: relative improvement of long-term setting. Right: visualization study of PEMS-08.
-As shown in Figure 3, we present the relative performance improvement of the advanced STID model with our ST-TTC method, and give a test set prediction visualization case on the PEMS-08 dataset (see Figure 9 in the Appendix for more examples).
+As shown in Figure 3, we present the relative performance improvement of the advanced STID model with our ST-TTC method, and give a test set prediction visualization case on the PEMS-08 dataset (see Figure F.1 in the Appendix for more examples).
 Our observations include: ❶ ST-TTC consistently improves long-term forecasting, even more than short-term (Table LABEL:tab:rq1_consistent_performance_12), likely due to more learnable information in longer windows.
@@
 We explore large-scale scenarios using the popular LargeST benchmark (comprising SD, GBA, GLA, and CA subsets).
-Figure 4 illustrates the 12-step prediction performance gains of the state-of-the-art efficient spatio-temporal model PatchSTG fang2025efficient with our ST-TTC , along with a comparison of inference time complexity (For full results, see Table 7 in Appendix).
+Figure 4 illustrates the 12-step prediction performance gains of the state-of-the-art efficient spatio-temporal model PatchSTG (Fang et al., 2025) with our ST-TTC , along with a comparison of inference time complexity (For full results, see Table E.2 in Appendix).
 We observe: ❶ Our ST-TTC consistently yields further performance improvements across all datasets, even surpassing the improvement of the second-best baseline over the PatchSTG on some datasets.
@@
 To illustrate the flexibility of ST-TTC in accommodating existing learning paradigms, we explore its integration with two training data-leveraging paradigms: OOD learning and Continual Learning.
-Figure 5: Relative improvement using our ST-TTC method in the OOD learning setting.
+Figure 5: Relative improvement using our ST-TTC in the OOD learning setting.
 OOD Learning Setting.
-Following prior work wang2024stone , we use the SD dataset to simulate spatio-temporal shift.
+Following prior work (Wang et al., 2024a), we use the SD dataset to simulate spatio-temporal shift.
 For the temporal dimension, we use 1-8/2019, 9-10/2019, and 11-12/2020 for training, validation, and testing, respectively.
 For the spatial dimension, we randomly mask 10% of nodes in the test set and consider three proportions of new nodes (10% / 15% / 20%) relative to the training node to mimic varying degrees of shift.
-In Figure 5, we present the 12-step average prediction performance gains of the advanced OOD learning model STONE with our ST-TTC , evaluated on all nodes and new nodes to demonstrate generalizability and scalability (Full results in Table 8).
+In Figure 5, we present the 12-step average prediction performance gains of the advanced OOD learning model STONE with our ST-TTC , evaluated on all nodes and new nodes to demonstrate generalizability and scalability (Full results in Table E.3).
 We observe that: ❶ The STONE model with ST-TTC consistently achieves performance benefits, significantly outperforming all previous settings, indicating that existing OOD models are still insufficient for true OOD generalization, while our method is highly effective.
@@
 Continual Learning Setting.
-Following prior work chen2025eac , we used multi-period streaming spatio-temporal data to examine our ST-TTC ’s integration with continual learning method.
+Following prior work (Chen and Liang, 2025), we used multi-period streaming spatio-temporal data to examine our ST-TTC ’s integration with continual learning method.
 Table 4 shows the improved 12-step forecasting of advanced continual learning models EAC and STKEC with our ST-TTC .
 We observed: ❶ Consistent performance gains for both models across all datasets; STKEC with ST-TTC even achieved comparable performance to best model EAC.
-(2) Energy-Stream achieves significant improvement over other datasets, as the ST-TTC effectively learns and calibrates temporal changes, as shown by the frequency analysis (drastic shift changes) in Figure 6.
+❷ Energy-Stream achieves significant improvement over other datasets, as the ST-TTC effectively learns and calibrates temporal changes, as shown by the frequency analysis (drastic shift changes) in Figure 6.
 Table 4: Performance comparison in continual learning setting.
@@
 MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
-EAC $24.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $38.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ $31.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $14.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ $20.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $5.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $50.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.60}}$
-$23.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $37.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ $31.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $14.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $23.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ $20.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ $3.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ $3.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.00}}$ $39.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}$
-$\Delta$downarrow 2.5% $\downarrow$ 1.9% $\downarrow$ 1.1% $\downarrow$ 1.4% $\downarrow$ 1.2% $\downarrow$ 1.4% $\downarrow$ 32.6% $\downarrow$ 27.8% $\downarrow$ 21.6%
-STKEC $25.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.05}}$ $40.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $33.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.64}}$ $16.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $26.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $22.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $5.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $52.40{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$
-$24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $39.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $31.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $16.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $21.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $4.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $43.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$
-$\Delta$downarrow 4.6% $\downarrow$ 2.7% $\downarrow$ 4.7% $\downarrow$ 1.2% $\downarrow$ 1.3% $\downarrow$ 2.0% $\downarrow$ 29.2% $\downarrow$ 25.2% $\downarrow$ 17.5%
+EAC $24.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ $38.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.31}}$ $31.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $14.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$ $20.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $5.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $5.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $50.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.60}}$
+$23.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $37.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ $31.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $14.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $23.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$ $20.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $3.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$ $3.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.00}}$ $39.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.41}}$
+$\Delta$ $\downarrow$ 2.5% $\downarrow$ 1.9% $\downarrow$ 1.1% $\downarrow$ 1.4% $\downarrow$ 1.2% $\downarrow$ 1.4% $\downarrow$ 32.6% $\downarrow$ 27.8% $\downarrow$ 21.6%
+STKEC $25.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.05}}$ $40.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.13}}$ $33.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.64}}$ $16.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.04}}$ $6.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}} 22$$2.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $5.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $5.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $52.40{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.10}}$
+$24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $39.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $31.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.04}}$ $16.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ $21.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $4.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $43.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.90}}$
+$\Delta$ $\downarrow$ 4.6% $\downarrow$ 2.7% $\downarrow$ 4.7% $\downarrow$ 1.2% $\downarrow$ 1.3% $\downarrow$ 2.0% $\downarrow$ 29.2% $\downarrow$ 25.2% $\downarrow$ 17.5%
 Figure 6: Energy-Stream’s Shift.
 5.5 Mechanism & Robustness Study (RQ4)
-Figure 7: Left: Strategy comparison. Middle: Effect of $lr$ w.r.t $m$ Right: Effect of $n$ w.r.t $s$.
+Figure 7: Left: Strategy comparison. Middle: Effect of lrlr w.r.t $m$ . Right: Effect of $n$ w.r.t $s$.
 We follow the OOD setup (challenging setting with 20% new nodes) to evaluate our ST-TTC .
@@
 We compare different strategies: 1) simple nonlinear time domain calibration (Time), 2) learning only phase or amplitude modulation factors (Pha.
-/ Amp.), 3) node-share modeling (Node), and (4) random selection or retrieval of the most similar samples (Rand.
+/ Amp.), 3) node-share modeling (Node), and 4) random selection or retrieval of the most similar samples (Rand.
 / Sim.).
@@
 In future work, we aim to explore how to enhance the internal computational capacity of spatio-temporal foundation models during test time.
-Supplementary material
-Learning with Calibration: Exploring
-Test-Time Computing of Spatio-Temporal Forecasting
-[…]
-[…]
+Acknowledgment
+The first author would like to thank all the anonymous reviewers for their valuable comments and high recognition.
+Although he has almost lost his passion for this field and is ready to leave, he hopes that this study can still bring something new to the ST community.
+This work is mainly supported by the National Natural Science Foundation of China (No.
+62402414).
+This work is also supported by the Huawei Co., Ltd (No.
+TC20241023027), the Guangzhou-HKUST(GZ) Joint Funding Program (No.
+2024A03J0620), Guangzhou Municipal Science and Technology Project (No.
+2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No.
+2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.
+2023B1212010007).
+Appendix
 Appendix A More Related Work
@@
 Many recent forecasting models leverage spectral (Fourier or wavelet) representations to capture periodic or multiscale patterns in spatio-temporal data.
-For example, PastNet [79] integrates a Fourier-domain convolutional operator to embed physical inductive biases, achieving state-of-the-art results in weather and traffic prediction.
-FourierGNN [85] builds a learnable Fourier-graph operator that conducts graph convolutions in the frequency domain, reducing convolutional complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$.
-Wavelet-based methods like WDNO [25] perform diffusion modeling in the wavelet domain to capture abrupt spatio-temporal changes and multi-resolution features.
-In the pure time-series setting, approaches such as FITS [84] interpolate in the complex Fourier domain and discard negligible high-frequency components to maintain accuracy with very few parameters, and TimeKAN [26] explicitly decomposes multivariate series into multiple frequency bands using Kolmogorov–Arnold networks.
+For example, PastNet (Wu et al., 2024) integrates a Fourier-domain convolutional operator to embed physical inductive biases, achieving state-of-the-art results in weather and traffic prediction.
+FourierGNN (Yi et al., 2023) builds a learnable Fourier-graph operator that conducts graph convolutions in the frequency domain, reducing convolutional complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$.
+Wavelet-based methods like WDNO (Hu et al., 2025) perform diffusion modeling in the wavelet domain to capture abrupt spatio-temporal changes and multi-resolution features.
+In the pure time-series setting, approaches such as FITS (Xu et al., 2024b) interpolate in the complex Fourier domain and discard negligible high-frequency components to maintain accuracy with very few parameters, and TimeKAN (Huang et al., 2025) explicitly decomposes multivariate series into multiple frequency bands using Kolmogorov–Arnold networks.
 These works demonstrate how frequency-domain learning can improve forecasting efficiency and accuracy by isolating dominant spectral components.
@@
 Recently, deep-learning approaches have been proposed to handle nonstationarity in an online fashion.
-For instance, FSNet [55] implements a complementary “fast and slow” learning system: a fast-adapting component for sudden pattern changes and a slow memory component for repeating trends.
-OneNet [78] runs two parallel neural forecasters (one modeling temporal dependencies, one modeling cross-variable dependencies) and uses reinforcement learning to dynamically weight their predictions under concept drift.
+For instance, FSNet (Pham et al., 2023) implements a complementary “fast and slow” learning system: a fast-adapting component for sudden pattern changes and a slow memory component for repeating trends.
+OneNet (Wen et al., 2023) runs two parallel neural forecasters (one modeling temporal dependencies, one modeling cross-variable dependencies) and uses reinforcement learning to dynamically weight their predictions under concept drift.
 These methods continuously update model parameters or ensemble weights as new data arrive.
-A recent study [35] pointed out the information leakage problem of previous online time series prediction methods, where the model makes predictions and then evaluates them based on the historical time steps that have been back-propagated for parameter updates.
+A recent study (Lau et al., 2025) pointed out the information leakage problem of previous online time series prediction methods, where the model makes predictions and then evaluates them based on the historical time steps that have been back-propagated for parameter updates.
 By redefining the setting to focus on predicting unknown future steps and evaluating unobserved data points, they propose a two-stream framework for online prediction, DSOF, which is conceptually similar to previous methods, generating predictions in a coarse-to-fine manner through a teacher-student model.
@@
 Entropy minimization methods adjust a trained model to increase prediction confidence on unlabeled test data.
-For example, [71] propose TENT, which adapts model parameters by minimizing the entropy of its predictions on each test batch and updating batch-normalization layers online.
+For example, (Wang et al., 2021) propose TENT, which adapts model parameters by minimizing the entropy of its predictions on each test batch and updating batch-normalization layers online.
 Feature alignment methods recalibrate feature distributions using test inputs; for instance, adaptive batch-normalization techniques re-estimate BN statistics on the target data to align feature distributions without labels.
 Self-supervised adaptation uses auxiliary tasks on the test data to refine the model.
-Test-Time Training [64, 18, 24, 76] converts each test input into a self-supervised learning problem (e.g.
+Test-Time Training (Gandelsman et al., 2022; Hardt and Sun, 2024; Sun et al., 2020; Wang et al., 2025a) converts each test input into a self-supervised learning problem (e.g.
 predicting image rotations) and updates model parameters before making a prediction.
-Similarly, SHOT [43] freezes the source classifier and updates the feature extractor on unlabeled target data using pseudo-labeling and information maximization.
+Similarly, SHOT (Liang et al., 2020) freezes the source classifier and updates the feature extractor on unlabeled target data using pseudo-labeling and information maximization.
 Each of these paradigms improves generalization under distribution shift without access to target labels.
@@
 Appendix B Theoretical Analysis
-B.1 Output Perturbation Bound
-Theorem 1 (Output Perturbation Bound).
+B.1 Approximate Bound on Output Perturbation
+Theorem 1 (Approximate Bound on Output Perturbation).
 Let $Y\in\mathbb{C}^{B\times N\times M}$ be the original frequency-domain representation of the backbone’s prediction $y\in\mathbb{R}^{B\times N\times T}$, and $y^{\prime}\in\mathbb{R}^{B\times N\times T}$ be the calibrated output.
@@
 $\|\Delta Y\|_{2}^{2}=\sum_{g=1}^{G}\sum_{f\in\text{Group }g}|\Delta Y_{g,f}|^{2}\leq\sum_{g=1}^{G}(\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2})\sum_{f\in\text{Group }g}|Y_{g,f}|^{2}=(\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2})\|Y\|_{2}^{2}.$
-By Parseval’s theorem [53], $\|y^{\prime}-y\|_{2}=\|\Delta Y\|_{2}$, hence:
+By Parseval’s theorem (Parseval, 1806), $\|y^{\prime}-y\|_{2}=\|\Delta Y\|_{2}$, hence:
 $\|y^{\prime}-y\|_{2}\leq\sqrt{\epsilon_{\alpha}^{2}+\epsilon_{\phi}^{2}}\|Y\|_{2}\leq(\epsilon_{\alpha}+\epsilon_{\phi})\|Y\|_{2}.$
 Remark 1.
-This theorem guarantees that the calibration-induced perturbation is linearly bounded by the modulation parameters $\epsilon_{\alpha},\epsilon_{\phi}$.
+This theorem guarantees that the calibration-induced perturbation is sub-linearly bounded by the modulation parameters $\epsilon_{\alpha},\epsilon_{\phi}$.
 By constraining these parameters (e.g., via regularization during test-time adaptation), SD-Calibrator ensures the calibrated output does not deviate excessively from the original prediction, thereby avoiding overfitting to transient noise.
@@
 $\|\nabla_{\lambda}L_{k}(\lambda_{a})-\nabla_{\lambda}L_{k}(\lambda_{b})\|_{2}\leq L_{c}\|\lambda_{a}-\lambda_{b}\|_{2}$
-According to the descent Lemma [48], this implies:
+According to the descent Lemma (Nesterov, 2013), this implies:
 $L_{k}(\lambda_{b})\leq L_{k}(\lambda_{a})+\langle\nabla_{\lambda}L_{k}(\lambda_{a}),\lambda_{b}-\lambda_{a}\rangle+\frac{L_{c}}{2}\|\lambda_{b}-\lambda_{a}\|_{2}^{2}$
@@
 Algorithm Workflow. We summarize the algorithm workflow of Section 4.1 in Algorithm 1.
+Algorithm Pseudo-code. We further present Algorithm 1 in the form of pytorch pseudo code in Listing LABEL:pytorch_sdc for easy understanding.
+C.2 Lightning Gradient Update
+Algorithm Pseudo-code. We summarize the algorithm workflow of Section 4.2 in Algorithm 2.
+Algorithm Workflow. We further present Algorithm 2 in the form of pytorch pseudo code in Listing LABEL:pytorch_fgu for easy understanding.
 Algorithm 1 Spectral Domain Calibrator
-0:  Pre-trained backbone $f_{\theta}$, Test input $x$, Horizon length $T$, Number of nodes $N$, Groups $G$
+0: Pre-trained backbone $f_{\theta}$, Test input $x$, Horizon length $T$, Number of nodes NN, Groups $G$
 0: Calibrated output $\hat{y}^{\mathrm{cal}}$
 1: Get the backbone predictions: $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N\times T}$
-2:  Compute $M\leftarrow\tfrac{T}{2}+1$vartriangleright I: Spatial-aware Decomposition
-3:  Apply real-to-complex FFT along time dimension for each node: $Y_{f}\;\leftarrow\;\mathrm{rFFT}(\hat{y})\;\in\;\mathbb{C}^{N\times M}$
-4:  Decompose: $A\leftarrow|Y_{f}|,\quad P\leftarrow\angle Y_{f}$vartriangleright II: Group-wise Modulation
+[…]
+3: Apply real‐to‐complex FFT along time dimension for each node: $Y_{f}\;\leftarrow\;\mathrm{rFFT}(\hat{y})\;\in\;\mathbb{C}^{N\times M}$
+[…]
 5:for $g=1,\dots,G$ do
@@
 12:return $\hat{y}^{\mathrm{cal}}$
-Algorithm Pseudo-code. We further present Algorithm 1 in the form of pytorch pseudo code in Algorithm 2 for easy understanding.
-Algorithm 2 PyTorch-style pseudocode: SD-Calibrator Class
-{minted}
-python class SD_Calibrator(nn.Module): """ Spectral Domain Calibrator with Phase-Amplitude Modulation """
-def __init__(self, num_nodes, freq_bins, groups=4): """ Args: num_nodes: number of spatial nodes (N) freq_bins: number of frequency bins (M = T // 2 + 1) groups: number of frequency groups (G) """ super().__init__() self.groups = groups self.group_size = freq_bins // groups
-# Learnable offsets for amplitude and phase: (G, N, 1) self.lambda_amp = nn.Parameter( torch.zeros(groups, num_nodes, 1) ) self.lambda_phi = nn.Parameter( torch.zeros(groups, num_nodes, 1) )
-def forward(self, y_pred): """ Args: y_pred: prediction from backbone, shape (B, 1, N, T) B defaults to 1, because only one sample can be tested Returns: calibrated prediction, shape (B, 1, N, T) """ B, _, N, T = y_pred.shape y = y_pred[:, 0] # (B, N, T)
-Yf = torch.fft.rfft(y, dim=-1) # (B, N, M) A = torch.abs(Yf) P = torch.angle(Yf)
-Yf_corr = torch.zeros_like(Yf) for g in range(self.groups): start = g * self.group_size if g == self.groups - 1 end = T // 2 + 1 else end = (g + 1) * self.group_size
-lam_a = self.lambda_amp[g].unsqueeze(0) # (1, N, 1) lam_p = self.lambda_phi[g].unsqueeze(0)
-A_g = A[:, :, start:end] * (1 + lam_a) P_g = P[:, :, start:end] + lam_p
+Listing 1: PyTorch-style pseudocode: SD-Calibrator Class
+class SD_Calibrator($n$.Module):
+Spectral Domain Calibrator with Phase-Amplitude Modulation
+def __init__(self, num_nodes, freq_bins, groups=4):
+Args:
+num_nodes: number of spatial nodes (N)
+freq_bins: number of frequency bins (M = T // 2 + 1)
+groups: number of frequency groups (G)
+super().__init__()
+self.groups = groups
+self.group_size = freq_bins // groups
+# Learnable offsets for amplitude and phase: (G, N, 1)
+self.lambda_amp = $n$.Parameter(
+torch.zeros(groups, num_nodes, 1)
+self.lambda_phi = $n$.Parameter(
+torch.zeros(groups, num_nodes, 1)
+def forward(self, y_pred):
+Args:
+y_pred: prediction from backbone, shape (B, 1, N, T)
+B defaults to 1, because only one sample can be tested
+Returns:
+calibrated prediction, shape (B, 1, N, T)
+B, _, N, T = y_pred.shape
+y = y_pred[:, 0] # (B, N, T)
+Yf = torch.fft.rfft(y, dim=-1) # (B, N, M)
+A = torch.abs(Yf)
+P = torch.angle(Yf)
+Yf_corr = torch.zeros_like(Yf)
+for g in range(self.groups):
+start = g * self.group_size
+if g == self.groups - 1
+end = T // 2 + 1
+else
+end = (g + 1) * self.group_size
+lam_a = self.lambda_amp[g].unsqueeze(0) # (1, N, 1)
+lam_p = self.lambda_phi[g].unsqueeze(0)
+A_g = A[:, :, start:end] * (1 + lam_a)
+P_g = P[:, :, start:end] + lam_p
 Yf_corr[:, :, start:end] = A_g * torch.exp(1j * P_g)
-y_time = torch.fft.irfft(Yf_corr, n=T, dim=-1) return y_time.unsqueeze(1) # (B, 1, N, T)
-C.2 Lightning Gradient Update
-Algorithm Pseudo-code. We summarize the algorithm workflow of Section 4.2 in Algorithm 3.
+y_time = torch.fft.irfft(Yf_corr, n=T, dim=-1)
+return y_time.unsqueeze(1) # (B, 1, N, T)
 Algorithm 2 Flash Gradient Update Mechanism
@@
 2:for each timestep $t=1,2,$... do
-3:     Receive $x_{t}$, compute default prediction $\hat{y}_{t}=f_{\theta}(x_{t})$vartriangleright I: Streaming Memory Queue
+[…]
 4:  Use Algorithm 1 to obtain the calibration results: $\hat{y}_{t}^{cal}=g_{\theta}(\hat{y}_{t}^{cal};\lambda)$
@@
 […]
-7:     if len($\mathcal{Q}$) $>T$ then
+7:  if len($\mathcal{Q}$) >T>T then
 8:   $(x_{o},y_{o})=\mathcal{Q}$.dequeue()
@@
 13:return $g_{\theta}$, $\{\hat{y}_{t}^{cal}\}_{t=1}^{B}$
-Algorithm Workflow. We further present Algorithm 3 in the form of pytorch pseudo code in Algorithm 4 for easy understanding.
-Algorithm 4 PyTorch-style pseudocode: Flash Gradient Update Function
-{minted}
-python def st_ttc_test(self, test_loader, node_num, T, groups): """ Flash Gradient Update with Streaming Memory Queue """ SDC = SD_Calibrator(node_num, T//2+1, groups).to(self.device) optimizer = torch.optim.Adam(SDC.parameters(), lr=1e-4) loss_fn = self._select_criterion() SMQ, preds = Queue(maxsize=T), []
-for x, y in test_loader: x, y = x.to(self.device), y.to(self.device) with torch.no_grad(): y_pred = self.model(x) y_corr = SDC(y_pred)
-# Use y_corr for inference y_corr = self.scaler.inverse_transform(y_corr) preds.append(y_corr.cpu().detach().numpy())
-SMQ.put((x, y)) if SMQ.full(): x_old, y_old = SMQ.get() with torch.no_grad(): y_pred_old = self.model(x_old)
-SDC.train() y_corr_old = SDC(y_pred_old) y_corr_old = self.scaler.inverse_transform(y_corr_old)
-loss = loss_fn(y_corr_old, y_old) loss.backward() optimizer.step() optimizer.zero_grad() SDC.eval() return preds
+Listing 2: PyTorch-style pseudocode: Flash Gradient Update Function
+def st_ttc_test(self, test_loader, node_num, T, groups):
+Flash Gradient Update with Streaming Memory Queue
+SDC = SD_Calibrator(node_num, T//2+1, groups).to(self.device)
+optimizer = torch.optim.Adam(SDC.parameters(), lr=1e-4)
+loss_fn = self._select_criterion()
+SMQ, preds = Queue(maxsize=T), []
+for x, y in test_loader:
+x, y = x.to(self.device), y.to(self.device)
+with torch.no_grad():
+y_pred = self.model(x)
+y_corr = SDC(y_pred)
+# Use y_corr for inference
+y_corr = self.scaler.inverse_transform(y_corr)
+preds.append(y_corr.cpu().detach().numpy())
+SMQ.put((x, y))
+if SMQ.full():
+x_old, y_old = SMQ.get()
+with torch.no_grad():
+y_pred_old = self.model(x_old)
+SDC.train()
+y_corr_old = SDC(y_pred_old)
+y_corr_old = self.scaler.inverse_transform(y_corr_old)
+loss = loss_fn(y_corr_old, y_old)
+loss.backward()
+optimizer.step()
+optimizer.zero_grad()
+SDC.eval()
+return preds
 Appendix D Experimental Details
 D.1 Datasets Details
-Our experiments are carried out on 14 real-world datasets from diffrent domain. The statistics of these spatio-temporal datasets are shown in Table 5.
-Table 5: Summary of datasets used for our experiments.
+Our experiments are carried out on 14 real-world datasets from diffrent domain. The statistics of these spatio-temporal datasets are shown in Table D.1.
+We follow the conventional practice (Li et al., 2018) to define the graph topology for all spatio-temporal datasets except Know-Air.
+Specifically, we construct the adjacency matrix $A$ for each dataset using a threshold Gaussian kernel, defined as follows:
+$A_{[ij]}=\begin{cases}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)&\text{if }\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)\geq r\penalty 10000\ \text{and}\penalty 10000\ i\neq j\\ 0&\text{otherwise}\end{cases}$
+where $d_{ij}$ represents the distance between sensors $i$ and $j$, $\sigma$ is the standard deviation of all distances, and $r$ is the threshold.
+We follow the recommended parameter settings in all corresponding papers.
+For the KnowAir dataset, we follow the original paper (Wang et al., 2020) and calculate the correlation between nodes to construct the adjacency matrix.
+Intuitively, most aerosol pollutants are distributed within a certain range above the ground.
+In addition, the mountains along the two cities will hinder the transmission of pollutants to the PM2.5 direction.
+Based on these intuitions, we constrain the weights in the adjacency matrix by the following formula:
+$A_{[ij]}=H(d_{\theta}-d_{ij})\cdot H(m_{\theta}-m_{ij}),\quad\text{where}$
+[…]
+where $\rho_{i}$ is the location (latitude, longitude) of node $i$, $h(\rho)$ is the height of location $\rho$, and $||\cdot||$ is the L2-norm of the vector.
+$H(\cdot)$ is the Heaviside step function, where $H(x)=1$ if and only if $x>0$.
+$d_{\theta}$ and $m_{\theta}$ are the distance and altitude thresholds, respectively.
+Specifically, we also set the distance threshold $d_{\theta}=300$ km and the altitude threshold $m_{\theta}=1200$ meters.
+Table D.1: Summary of datasets used for our experiments.
 Degree: the average degree of each node.
@@
 Source Dataset Nodes Time Range Frames Sampling Rate Data Points
- [63] PEMS03 358 09/01/2018 – 11/30/2018 26,208 5 minutes 9.38M
+(Song et al., 2020) PEMS03 358 09/01/2018 – 11/30/2018 26,208 5 minutes 9.38M
 PEMS04 307 01/01/2018 – 02/28/2018 16,992 5 minutes 5.22M
@@
 PEMS08 170 07/01/2016 – 08/31/2016 17,856 5 minutes 3.04M
- [37] UrbanEV 275 09/01/2022 – 02/28/2023 4344 1 hour 1.19M
- [77] Know-Air 184 01/01/2015 – 12/31/2018 11688 3 hours 2.15M
- [38] METR-LA 207 03/01/2012 – 06/27/2012 34,272 5 minutes 7.09M
-LargeST [46] CA 8,600 01/01/2019 – 12/31/2019 35,040 15 minutes 30.13M
+(Li et al., 2025) UrbanEV 275 09/01/2022 – 02/28/2023 4344 1 hour 1.19M
+(Wang et al., 2020) Know-Air 184 01/01/2015 – 12/31/2018 11688 3 hours 2.15M
+(Li et al., 2018) METR-LA 207 03/01/2012 – 06/27/2012 34,272 5 minutes 7.09M
+LargeST (Liu et al., 2023b) CA 8,600 01/01/2019 – 12/31/2019 35,040 15 minutes 30.13M
 GLA 3,834 01/01/2019 – 12/31/2019 35,040 15 minutes 13.43M
@@
 SD 716 01/01/2019 – 12/31/2020 70,080 15 minutes 5.02M
- [8] Air-Stream 1087 $\rightarrow$ 1154 $\rightarrow$ 1193 $\rightarrow$ 1202 01/01/2016 - 12/31/2019 34065 1 hour 15.79M
-PEMS-Stream 655 $\rightarrow$ 715 $\rightarrow$ 786 $\rightarrow$ 822 $\rightarrow$ 834 $\rightarrow$ 850 $\rightarrow$ 871 07/10/2011 - 09/08/2017 61,992 5 minutes 34.30M
-Energy-Stream 103 $\rightarrow$ 113 $\rightarrow$ 122 $\rightarrow$ 134 Unknown (245 days) 34,560 10 minutes 1.63M
-We follow the conventional practice [38] to define the graph topology for all spatio-temporal datasets except Know-Air.
-Specifically, we construct the adjacency matrix $A$ for each dataset using a threshold Gaussian kernel, defined as follows:
-$A_{[ij]}=\begin{cases}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)&\text{if}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)\geq r~{}\text{and}~{}i\neq j\\ 0&\text{otherwise}\end{cases}$
-where $d_{ij}$ represents the distance between sensors $i$ and $j$, $\sigma$ is the standard deviation of all distances, and $r$ is the threshold.
-We follow the recommended parameter settings in all corresponding papers.
-For the KnowAir dataset, we follow the original paper [77] and calculate the correlation between nodes to construct the adjacency matrix.
-Intuitively, most aerosol pollutants are distributed within a certain range above the ground.
-In addition, the mountains along the two cities will hinder the transmission of pollutants to the PM2.5 direction.
-Based on these intuitions, we constrain the weights in the adjacency matrix by the following formula:
-$A_{[ij]}=H(d_{\theta}-d_{ij})\cdot H(m_{\theta}-m_{ij}),\quad\text{where}$
-$d_{ij}=||\rho_{i}-\rho_{j}||,\quad m_{ij}=\sup_{\lambda\in(0,1)}\{h(\lambda\rho_{i}+(1-\lambda)\rho_{j})-\max\{h(\rho_{i}),h(\rho_{j})\}\},$
-where $\rho_{i}$ is the location (latitude, longitude) of node $i$, $h(\rho)$ is the height of location $\rho$, and |||||$\cdot$||| | ⋅ | | is the L2-norm of the vector.
-$H(\cdot)$ is the Heaviside step function, where $H(x)=1$ if and only if $x>0$.
-$d_{\theta}$ and $m_{\theta}$ are the distance and altitude thresholds, respectively.
-Specifically, we also set the distance threshold $d_{\theta}=300$ km and the altitude threshold $m_{\theta}=1200$ meters.
+(Chen and Liang, 2025) Air-Stream
+1087 $\rightarrow$ 1154
+$\rightarrow$ 1193 $\rightarrow$ 1202
+01/01/2016 - 12/31/2019 34065 1 hour 15.79M
+PEMS-Stream
+655 $\rightarrow$ 715 $\rightarrow$ 786
+$\rightarrow$ 822 $\rightarrow$ 834 $\rightarrow$ 850
+$\rightarrow$ 871
+07/10/2011 - 09/08/2017 61,992 5 minutes 34.30M
+Energy-Stream
+103 $\rightarrow$ 113
+$\rightarrow$ 122 $\rightarrow$ 134
+Unknown (245 days) 34,560 10 minutes 1.63M
 D.2 Baseline Details
@@
 Classical Learning Methods for Spatio-Temporal Forecasting.
-  • STAEformer [45]: STAEformer is a spatial-temporal adaptive embedding transformer that makes vanilla transformer state-of-the-art for spatio-temporal forecasting.
+  • STAEformer (Liu et al., 2023a): STAEformer is a spatial-temporal adaptive embedding transformer that makes vanilla transformer state-of-the-art for spatio-temporal forecasting.
 It introduces a novel architecture to effectively capture the dynamic spatial and temporal dependencies in spatio-temporal data.
 https://github.com/XDZhelheim/STAEformer
-  • STTN [83]: STTN is a spatial-temporal transformer network designed for traffic flow forecasting.
+  • STTN (Xu et al., 2020): STTN is a spatial-temporal transformer network designed for traffic flow forecasting.
 It leverages dynamic directed spatial dependencies and long-range temporal dependencies to enhance the accuracy of long-term traffic predictions.
 https://github.com/xumingxingsjtu/STTN
-  • GWNet [80]: GWNet is a graph wavenet model for deep spatial-temporal graph modeling.
+  • GWNet (Wu et al., 2019): GWNet is a graph wavenet model for deep spatial-temporal graph modeling.
 It effectively captures the complex spatial and temporal patterns in spatio-temporal data using a combination of graph convolutional networks and dilated causal convolutions.
 https://github.com/nnzhan/Graph-WaveNet
-  • STGCN [87]: STGCN is a spatio-temporal graph convolutional network framework for traffic forecasting.
+  • STGCN (Yu et al., 2018): STGCN is a spatio-temporal graph convolutional network framework for traffic forecasting.
 It integrates graph convolutional networks with temporal convolutional networks to model the spatial and temporal dependencies in traffic data.
 https://github.com/hazdzz/stgcn
-  • STID [60]: STID is a simple yet effective baseline for spatio-temporal forecasting.
+  • STID (Shao et al., 2022b): STID is a simple yet effective baseline for spatio-temporal forecasting.
 It addresses the indistinguishability of samples in spatial and temporal dimensions by attaching spatial and temporal identity information, achieving competitive performance with concise and efficient models.
 https://github.com/GestaltCogTeam/STID
-  • ST-Norm [13]: ST-Norm is a method that applies spatial and temporal normalization for multi-variate time series forecasting.
+  • ST-Norm (Deng et al., 2021): ST-Norm is a method that applies spatial and temporal normalization for multi-variate time series forecasting.
 It enhances the performance of forecasting models by normalizing the spatial and temporal features of the data.
@@
 Efficient Learning Methods for Large-Scale Spatio-Temporal Forecasting.
-  • PatchSTG [15]: PatchSTG is an attention-based dynamic spatial modeling method that uses irregular spatial patching for efficient large-scale spatio-temporal forecasting.
+  • PatchSTG (Fang et al., 2025): PatchSTG is an attention-based dynamic spatial modeling method that uses irregular spatial patching for efficient large-scale spatio-temporal forecasting.
 It reduces computational complexity by segmenting large-scale inputs into balanced and non-overlapped patches, capturing local and global spatial dependencies effectively.
@@
 OOD Learning Methods for Spatio-Temporal Forecasting.
-  • STONE [66]: STONE is a state-of-the-art spatio-temporal OOD learning framework that effectively models spatial heterogeneity and generates temporal and spatial semantic graphs.
+  • STONE (Wang et al., 2024a): STONE is a state-of-the-art spatio-temporal OOD learning framework that effectively models spatial heterogeneity and generates temporal and spatial semantic graphs.
 It introduces a graph perturbation mechanism to enhance the model’s environmental modeling capability for better generalization.
@@
 Continual Learning Methods for Spatio-Temporal Forecasting.
-  • EAC [8]: EAC is a state-of-the-art method for exploring the rapid adaptation of models in the face of dynamic spatio-temporal graph changes during supervised finetuning.
+  • EAC (Chen and Liang, 2025): EAC is a state-of-the-art method for exploring the rapid adaptation of models in the face of dynamic spatio-temporal graph changes during supervised finetuning.
 It follows the principles of expand and compress to address the challenges of retraining models over new data and catastrophic forgetting.
 https://github.com/Onedean/EAC
-  • STKEC [68]: STKEC is a continual learning framework for traffic flow prediction on expanding traffic networks.
+  • STKEC (Wang et al., 2023a): STKEC is a continual learning framework for traffic flow prediction on expanding traffic networks.
 It introduces a pattern bank to store representative network patterns and employs a pattern expansion mechanism to incorporate new patterns from evolving networks without requiring historical graph data.
@@
 Popular test-time training methods
-  • TTT-MAE [18]: TTT-MAE is a test-time training method that uses masked autoencoders to adjust the model during inference.
+  • TTT-MAE (Gandelsman et al., 2022): TTT-MAE is a test-time training method that uses masked autoencoders to adjust the model during inference.
 It helps improve the performance of the model on unseen data by effectively utilizing test-time information.
@@
 https://github.com/Rima-ag/TTT-MAE
-  • TENT [71]: TENT is a method for adjusting the model at test time by normalizing the activation function to reduce the offset between the training distribution and the test distribution.
+  • TENT (Wang et al., 2021): TENT is a method for adjusting the model at test time by normalizing the activation function to reduce the offset between the training distribution and the test distribution.
 It enhances the generalization ability of the model without retraining on labeled test data.
@@
 Classical online time series forecasting methods
-  • OnlineTCN [101]: OnlineTCN is an online learning method based on a time convolutional network.
+  • OnlineTCN (Zinkevich, 2003): OnlineTCN is an online learning method based on a time convolutional network.
 It can adapt to new data sequentially and is very suitable for real-time prediction applications where data arrives continuously.
 https://github.com/locuslab/TCN
-  • FSNet [55]: FSNet proposes a fast and slow learning network for online time series prediction that can handle both sudden changes and repeated patterns.
+  • FSNet (Pham et al., 2023): FSNet proposes a fast and slow learning network for online time series prediction that can handle both sudden changes and repeated patterns.
 In particular, FSNet improves on a slowly learning backbone by dynamically balancing fast adaptation to recent changes and retrieval of similar old knowledge.
@@
 https://github.com/salesforce/fsnet
-  • OneNet [78]: OneNet dynamically updates and combines two models, one focusing on modeling dependencies across time dimensions and the other focusing on cross-variable dependencies.
+  • OneNet (Wen et al., 2023): OneNet dynamically updates and combines two models, one focusing on modeling dependencies across time dimensions and the other focusing on cross-variable dependencies.
 The approach integrates reinforcement learning-based methods into a traditional online convex programming framework, allowing the two models to be linearly combined with dynamically adjusted weights, thereby addressing the main drawback of classical online prediction methods that are slow to adapt to concept drift.
@@
 Advanced spatio-temporal forecasting methods using test information.
-  • CompFormer [97]: CompFormer proposes a test-time compensated representation learning framework, including a spatiotemporal decomposed database and a multi-head spatial transformer model.
+  • CompFormer (Zhang et al., 2023c): CompFormer proposes a test-time compensated representation learning framework, including a spatiotemporal decomposed database and a multi-head spatial transformer model.
 The former component explicitly separates all training data along the time dimension according to periodic features, while the latter component establishes connections between recent observations and historical sequences in the database through a spatial attention matrix.
 This enables it to transfer robust features to overcome abnormal events
-  • DOST [70]: DOST proposes a novel online continuous learning framework tailored to the characteristics of spatiotemporal data.
+  • DOST (Wang et al., 2024c): DOST proposes a novel online continuous learning framework tailored to the characteristics of spatiotemporal data.
 DOSTadopts an adaptive spatiotemporal network equipped with variable independent adapters to dynamically address the unique distribution changes of each urban location.
@@
 Metrics Detail. We use different metrics such as MAE, RMSE, and MAPE. Formally, these metrics are formulated as following:
-$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,~{}~{}~{}\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},~{}~{}~{}\text{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|$
+$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|$
 where $n$ represents the indices of all observed samples, $y_{i}$ denotes the $i$-th actual sample, and $\hat{y_{i}}$ is the corresponding prediction.
@@
 All experiments are conducted on a Linux server equipped with a 1 × AMD EPYC 7763 128-Core Processor CPU (256GB memory) and 4 × NVIDIA RTX A6000 (48GB memory) GPUs.
-To carry out benchmark testing experiments, all baselines are set to run for a duration of 100$\sim$150 epochs by default (depends on the corresponding paper), with specific timings contingent upon the method with early stop mechanism.
+To carry out benchmark testing experiments, all baselines are set to run for a duration of $0\sim1$50 epochs by default (depends on the corresponding paper), with specific timings contingent upon the method with early stop mechanism.
 The number of early stopping steps is set to 10.
@@
 E.1 Complete Results Table
-We provide complete information of the experimental tables in the main text as Table LABEL:tab:rq2_fewshot_1278
-Table 6: Performance comparison of different models w/ and w/o ST-TTC in the few-shot scenario.
+We provide complete information of the experimental tables in the main text as Table LABEL:tab:rq2_fewshot_12, E.2, E.3
+Table E.1: Performance comparison of different models w/ and w/o ST-TTC in the few-shot scenario.
 Models Transformer-based Graph-based MLP-based
-STAEformer [45] STTN [83] GWNet [80] STGCN [87] STID [60] ST-Norm [13]
+STAEformer (Liu et al., 2023a) STTN (Xu et al., 2020) GWNet (Wu et al., 2019) STGCN (Yu et al., 2018) STID (Shao et al., 2022b) ST-Norm (Deng et al., 2021)
 w/ ST-TTC
-PEMS-03 MAE $23.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$ 22.96±0.93 $21.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.93}}$ 21.15±0.92 $21.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.98}}$ 21.43±0.92 $21.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 21.28±0.54 $21.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}$ 21.57±0.24 $21.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ 20.74±0.28
-RMSE $37.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ 37.10±1.59 $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.59}}$ 33.87±1.53 $34.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.44}}$ 34.28±1.40 $34.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.92}}$ 34.08±0.89 $35.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ 35.18±0.52 $33.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.45}}$ 33.07±0.35
-MAPE(%) $21.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.21}}$ 21.23±1.03 $21.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.35}}$ 21.13±1.28 $21.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.14}}$ 19.70±0.57 $22.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ 21.59±0.77 $21.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.86}}$ 21.27±0.73 $22.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.13}}$ 22.04±2.22
-PEMS-04 MAE $35.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.25}}$ 34.57±4.08 $29.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.37}}$ 29.44±0.31 $33.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.86}}$ 32.87±1.95 $29.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ 29.66±0.83 $29.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 29.49±0.65 $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ 30.31±0.14
-RMSE $50.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.86}}$ 50.23±4.59 $44.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.75}}$ 43.89±0.81 $50.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.56}}$ 49.73±2.73 $45.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.26}}$ 45.47±1.28 $44.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ 44.62±1.10 $45.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 45.33±0.49
-MAPE(%) $23.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.28}}$ 23.40±3.26 $23.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.27}}$ 22.54±0.71 $22.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.52}}$ 22.43±3.17 $20.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ 20.72±1.19 $22.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.77}}$ 22.70±1.72 $21.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 21.72±0.60
-PEMS-07 MAE $30.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.47}}$ 29.70±0.39 $31.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ 31.22±0.69 $33.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.65}}$ 32.82±0.63 $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.72}}$ 31.88±0.77 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.00}}$ 30.91±1.05 $31.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 30.50±0.05
-RMSE $47.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ 46.89±0.76 $46.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$ 45.96±0.93 $49.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 49.36±0.57 $48.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ 47.61±0.05 $47.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.82}}$ 46.71±0.97 $47.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.43}}$ 46.54±0.42
-MAPE(%) $13.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 13.53±0.23 $14.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}$ 14.45±0.13 $15.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.70}}$ 14.74±0.58 $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}$ 16.28±0.99 $15.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.15}}$ 15.03±1.20 $14.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 14.18±0.46
-PEMS-08 MAE $36.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.31}}$ 35.47±6.03 $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.42}}$ 23.81±0.41 $26.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.85}}$ 25.94±0.97 $25.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.25}}$ 25.31±0.21 $24.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 23.75±0.28 $24.34{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 24.06±0.07
-RMSE $54.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±10.46}}$ 52.43±8.37 $36.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.45}}$ 36.61±0.50 $40.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.95}}$ 40.51±1.11 $38.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ 37.80±0.10 $37.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.77}}$ 37.36±0.77 $37.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ 37.20±0.20
-MAPE(%) $27.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±9.55}}$ 26.19±8.29 $18.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ 17.65±0.54 $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$ 16.52±1.25 $20.16{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.91}}$ 17.84±0.88 $15.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ 14.43±0.20 $15.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 14.99±0.29
-KnowAir MAE $18.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 18.16±0.35 $20.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ 19.85±0.23 $19.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.32}}$ 19.04±0.29 $20.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ 20.09±0.25 $22.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.20}}$ 21.72±0.94 $21.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.39}}$ 20.92±0.36
-RMSE $27.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 26.97±0.08 $28.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ 28.51±0.46 $27.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ 27.58±0.15 $29.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.20}}$ 28.65±0.21 $30.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.03}}$ 29.69±0.83 $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ 28.86±0.41
-MAPE(%) $72.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.17}}$ 70.14±4.89 $85.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.82}}$ 81.21±1.63 $80.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.43}}$ 78.49±5.40 $84.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.40}}$ 82.13±2.18 $102.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.60}}$ 95.69±5.08 $95.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.26}}$ 91.11±1.80
-UrbanEV MAE $3.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ 3.36±0.07 $4.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 4.05±0.06 $3.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ 3.88±0.09 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 4.10±0.10 $3.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 3.24±0.05 $4.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 4.75±0.11
-RMSE $6.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.21}}$ 6.05±0.16 $6.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ 6.71±0.05 $6.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ 6.59±0.11 $6.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ 6.61±0.17 $5.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ 5.42±0.08 $8.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ 8.36±0.27
-MAPE(%) $31.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.19}}$ 31.33±0.47 $38.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.43}}$ 37.32±1.33 $35.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.24}}$ 34.95±1.10 $40.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.48}}$ 39.57±1.52 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.59}}$ 30.75±0.50 $43.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.28}}$ 42.26±1.25
-Table 7: PatchSTG with ST-TTC in LargeST Benchmark.
+PEMS-03 MAE $3.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.90}} 22$.96±0.93 $21.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.93}}$ 21.15±0.92 $21.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.98}}$ 21.43±0.92 $21.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.50}}$ 21.28±0.54 $21.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.24}}$ 21.57±0.24 $21.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.31}}$ 20.74±0.28
+RMSE $37.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.41}}$ 37.10±1.59 $4.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.59}} 33$.87±1.53 $34.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.44}}$ 34.28±1.40 $34.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.92}}$ 34.08±0.89 $35.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.53}}$ 35.18±0.52 $3.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.45}} 33$.07±0.35
+MAPE(%) $21.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.21}}$ 21.23±1.03 $21.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.35}}$ 21.13±1.28 $21.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.14}}$ 19.70±0.57 $22.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.76}}$ 21.59±0.77 $21.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.86}}$ 21.27±0.73 $2.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.13}} 22$.04±2.22
+PEMS-04 MAE $35.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.25}}$ 34.57±4.08 $29.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.37}}$ 29.44±0.31 $33.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.86}}$ 32.87±1.95 $29.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ 29.66±0.83 $29.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ 29.49±0.65 $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ 30.31±0.14
+RMSE $50.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.86}}$ 50.23±4.59 $44.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.75}}$ 43.89±0.81 $50.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.56}}$ 49.73±2.73 $45.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.26}}$ 45.47±1.28 $4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.13}} 44$.62±1.10 $45.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.50}}$ 45.33±0.49
+MAPE(%) $23.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.28}}$ 23.40±3.26 $3.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.27}} 22$.54±0.71 $2.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.52}} 22$.43±3.17 $20.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.19}}$ 20.72±1.19 $2.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.77}} 22$.70±1.72 $21.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ 21.72±0.60
+PEMS-07 MAE $30.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.47}}$ 29.70±0.39 $31.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.82}}$ 31.22±0.69 $33.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.65}}$ 32.82±0.63 $32.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.72}}$ 31.88±0.77 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.00}}$ 30.91±1.05 $31.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ 30.50±0.05
+RMSE $47.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ 46.89±0.76 $46.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.10}}$ 45.96±0.93 $49.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.59}}$ 49.36±0.57 $48.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ 47.61±0.05 $47.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.82}}$ 46.71±0.97 $47.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.43}}$ 46.54±0.42
+MAPE(%) $13.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ 13.53±0.23 $14.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$ 14.45±0.13 $15.04{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.70}}$ 14.74±0.58 $17.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.40}}$ 16.28±0.99 $15.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.15}}$ 15.03±1.20 $14.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ 14.18±0.46
+PEMS-08 MAE $36.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 7.31}}$ 35.47±6.03 $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.42}}$ 23.81±0.41 $26.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.85}}$ 25.94±0.97 $25.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.25}}$ 25.31±0.21 $24.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ 23.75±0.28 $24.34{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ 24.06±0.07
+RMSE $54.61{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 10.46}}$ 52.43±8.37 $36.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.45}}$ 36.61±0.50 $40.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.95}}$ 40.51±1.11 $38.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ 37.80±0.10 $37.62{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.77}}$ 37.36±0.77 $37.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ 37.20±0.20
+MAPE(%) $27.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 9.55}}$ 26.19±8.29 $18.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.52}}$ 17.65±0.54 $17.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.10}}$ 16.52±1.25 $20.16{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.91}}$ 17.84±0.88 $15.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.53}}$ 14.43±0.20 $15.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.59}}$ 14.99±0.29
+KnowAir MAE $18.48{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.50}}$ 18.16±0.35 $20.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}}$ 19.85±0.23 $19.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.32}}$ 19.04±0.29 $20.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$ 20.09±0.25 $22.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.20}}$ 21.72±0.94 $21.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.39}}$ 20.92±0.36
+RMSE $27.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ 26.97±0.08 $28.95{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$ 28.51±0.46 $27.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.13}}$ 27.58±0.15 $29.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.20}}$ 28.65±0.21 $30.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.03}}$ 29.69±0.83 $29.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$ 28.86±0.41
+MAPE(%) $72.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 7.17}}$ 70.14±4.89 $85.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.82}}$ 81.21±1.63 $80.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 5.43}}$ 78.49±5.40 $84.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.40}}$ 82.13±2.18 $102.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 6.60}}$ 95.69±5.08 $95.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.26}}$ 91.11±1.80
+UrbanEV MAE $39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}} 3.3$6±0.07 $4.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ 4.05±0.06 $3.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ 3.88±0.09 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ 4.10±0.10 $3.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ 3.24±0.05 $4.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ 4.75±0.11
+RMSE $6.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.21}}$ 6.05±0.16 $6.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ 6.71±0.05 $6.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$ 6.59±0.11 $74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.19}} 6.6$1±0.17 $5.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ 5.42±0.08 $8.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ 8.36±0.27
+MAPE(%) $31.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.19}}$ 31.33±0.47 $38.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.43}}$ 37.32±1.33 $35.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.24}}$ 34.95±1.10 $40.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.48}}$ 39.57±1.52 $31.42{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.59}}$ 30.75±0.50 $43.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.28}}$ 42.26±1.25
+Table E.2: PatchSTG with ST-TTC in LargeST Benchmark.
 Datasets Methods Horizon 3 Horizon 6 Horizon 12 Average
@@
 w/ ST-TTC 14.44 24.07 8.70 16.66 28.18 10.93 20.37 35.68 14.27 16.72 28.83 11.11
-$\Delta$downarrow 0.6% $\downarrow$ 1.1% $\downarrow$ 5.6% $\downarrow$ 1.2% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 3.1% $\downarrow$ 1.1% $\downarrow$ 1.5% $\downarrow$ 1.1%
+$\Delta$ $\downarrow$ 0.6% $\downarrow$ 1.1% $\downarrow$ 5.6% $\downarrow$ 1.2% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 3.1% $\downarrow$ 1.1% $\downarrow$ 1.5% $\downarrow$ 1.1%
 GBA PatchSTG 16.81 28.71 12.25 19.68 33.09 14.51 23.49 39.23 18.93 19.50 33.16 14.64
 w/ ST-TTC 16.65 28.31 12.12 19.30 32.40 14.35 22.96 38.33 18.30 19.16 32.49 14.48
-$\Delta$downarrow 1.0% $\downarrow$ 1.4% $\downarrow$ 1.1% $\downarrow$ 1.9% $\downarrow$ 2.1% $\downarrow$ 1.1% $\downarrow$ 2.3% $\downarrow$ 2.3% $\downarrow$ 3.3% $\downarrow$ 1.7% $\downarrow$ 2.0% $\downarrow$ 1.1%
+$\Delta$ $\downarrow$ 1.0% $\downarrow$ 1.4% $\downarrow$ 1.1% $\downarrow$ 1.9% $\downarrow$ 2.1% $\downarrow$ 1.1% $\downarrow$ 2.3% $\downarrow$ 2.3% $\downarrow$ 3.3% $\downarrow$ 1.7% $\downarrow$ 2.0% $\downarrow$ 1.1%
 GLA PatchSTG 15.84 26.34 9.27 19.06 31.85 11.30 23.32 39.64 14.60 18.96 32.33 11.44
 w/ ST-TTC 15.78 26.08 9.15 18.76 31.29 11.21 22.86 38.89 14.35 18.69 31.78 11.36
-$\Delta$downarrow 0.4% $\downarrow$ 1.0% $\downarrow$ 1.3% $\downarrow$ 1.6% $\downarrow$ 1.8% $\downarrow$ 0.8% $\downarrow$ 2.0% $\downarrow$ 1.9% $\downarrow$ 1.7% $\downarrow$ 1.4% $\downarrow$ 1.7% $\downarrow$ 0.7%
+$\Delta$ $\downarrow$ 0.4% $\downarrow$ 1.0% $\downarrow$ 1.3% $\downarrow$ 1.6% $\downarrow$ 1.8% $\downarrow$ 0.8% $\downarrow$ 2.0% $\downarrow$ 1.9% $\downarrow$ 1.7% $\downarrow$ 1.4% $\downarrow$ 1.7% $\downarrow$ 0.7%
 CA PatchSTG 14.69 24.82 10.51 17.41 29.43 12.83 21.20 36.13 16.00 17.35 29.79 12.79
 w/ ST-TTC 14.59 24.61 10.40 17.14 28.97 12.51 20.76 35.38 15.54 17.10 29.31 12.53
-$\Delta$downarrow 0.7% $\downarrow$ 0.8% $\downarrow$ 1.0% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 2.5% $\downarrow$ 2.1% $\downarrow$ 2.1% $\downarrow$ 2.9% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 2.0%
-Table 8: Performance of spatio-temporal shift dataset SD-ratio(%) on all nodes and unknown new nodes at different spatio-temporal shift levels.
+$\Delta$ $\downarrow$ 0.7% $\downarrow$ 0.8% $\downarrow$ 1.0% $\downarrow$ 1.6% $\downarrow$ 1.6% $\downarrow$ 2.5% $\downarrow$ 2.1% $\downarrow$ 2.1% $\downarrow$ 2.9% $\downarrow$ 1.4% $\downarrow$ 1.6% $\downarrow$ 2.0%
+Table E.3: Performance of spatio-temporal shift dataset SD-ratio(%) on all nodes and unknown new nodes at different spatio-temporal shift levels.
 Dataset Horizon Methods All Node New Node
 MAE RMSE MAPE(%) MAE RMSE MAPE(%)
-SD-10% 12 STONE $40.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.24}}$ $58.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.67}}$ $45.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±9.30}}$ $46.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.02}}$ $66.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±7.05}}$ $47.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±8.02}}$
-w/ ST-TTC $39.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.82}}$ $57.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.67}}$ $38.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.19}}$ $44.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.69}}$ $65.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±6.95}}$ $41.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±5.87}}$
-$\Delta$downarrow 3.3% $\downarrow$ 1.7% $\downarrow$ 17.0% $\downarrow$ 3.5% $\downarrow$ 2.9% $\downarrow$ 12.0%
+SD-10% 12 STONE $40.74{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.24}}$ $58.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.67}}$ $45.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 9.30}}$ $46.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 7.02}}$ $66.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 7.05}}$ $47.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 8.02}}$
+w/ ST-TTC $39.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.82}}$ $57.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.67}}$ $38.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 5.19}}$ $44.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 6.69}}$ $65.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 6.95}}$ $41.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 5.87}}$
+$\Delta$ $\downarrow$ 3.3% $\downarrow$ 1.7% $\downarrow$ 17.0% $\downarrow$ 3.5% $\downarrow$ 2.9% $\downarrow$ 12.0%
 Avg.
-STONE $30.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.19}}$ $42.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.38}}$ $39.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.03}}$ $32.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.77}}$ $46.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.87}}$ $40.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.30}}$
-w/ ST-TTC $28.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.04}}$ $41.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.16}}$ $28.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.32}}$ $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.63}}$ $44.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.32}}$ $30.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.54}}$
-$\Delta$downarrow 6.3% $\downarrow$ 3.1% $\downarrow$ 27.3% $\downarrow$ 6.5% $\downarrow$ 4.8% $\downarrow$ 23.2%
-SD-15% 12 STONE $35.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.46}}$ $52.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}$ $34.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.73}}$ $43.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.85}}$ $66.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.28}}$ $30.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.40}}$
-w/ ST-TTC $34.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $52.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.54}}$ $29.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $41.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ $63.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $27.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$
-$\Delta$downarrow 1.3% $\downarrow$ 1.2% $\downarrow$ 12.5% $\downarrow$ 3.9% $\downarrow$ 4.4% $\downarrow$ 9.1%
+STONE $30.18{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.19}}$ $42.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.38}}$ $39.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.03}}$ $32.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.77}}$ $46.68{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.87}}$ $40.12{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.30}}$
+w/ ST-TTC $28.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.04}}$ $41.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.16}}$ $28.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.32}}$ $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.63}}$ $44.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.32}}$ $30.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.54}}$
+$\Delta$ $\downarrow$ 6.3% $\downarrow$ 3.1% $\downarrow$ 27.3% $\downarrow$ 6.5% $\downarrow$ 4.8% $\downarrow$ 23.2%
+SD-15% 12 STONE $35.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.46}}$ $52.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.57}}$ $34.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.73}}$ $43.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.85}}$ $66.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.28}}$ $30.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.40}}$
+w/ ST-TTC $34.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $52.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.54}}$ $29.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.79}}$ $41.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.62}}$ $63.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $27.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$
+$\Delta$ $\downarrow$ 1.3% $\downarrow$ 1.2% $\downarrow$ 12.5% $\downarrow$ 3.9% $\downarrow$ 4.4% $\downarrow$ 9.1%
 Avg.
-STONE $28.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $41.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ $36.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.91}}$ $33.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.69}}$ $49.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ $29.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.00}}$
-w/ ST-TTC $26.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ $39.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $24.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.00}}$ $30.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $46.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ $22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}$
-$\Delta$downarrow 6.5% $\downarrow$ 3.4% $\downarrow$ 30.7% $\downarrow$ 7.7% $\downarrow$ 6.0% $\downarrow$ 24.4%
-SD-20% 12 STONE $36.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ $53.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $41.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.23}}$ $63.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.11}}$ $37.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.29}}$
-w/ ST-TTC $35.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $52.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ $30.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $40.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ $60.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.07}}$ $33.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.73}}$
-$\Delta$downarrow 2.9% $\downarrow$ 2.8% $\downarrow$ 10.6% $\downarrow$ 3.7% $\downarrow$ 3.8% $\downarrow$ 9.8%
+STONE $28.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $41.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$ $6.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.91}} 33$$3.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.69}}$ $49.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.97}}$ $29.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.00}}$
+w/ ST-TTC $26.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$ $39.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $24.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.00}}$ $30.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $46.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.52}}$ $22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.38}}$
+$\Delta$ $\downarrow$ 6.5% $\downarrow$ 3.4% $\downarrow$ 30.7% $\downarrow$ 7.7% $\downarrow$ 6.0% $\downarrow$ 24.4%
+SD-20% 12 STONE $36.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.76}}$ $53.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$ $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.08}}$ $41.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.23}}$ $63.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.11}}$ $37.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.29}}$
+w/ ST-TTC $35.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.08}}$ $52.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ $30.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.13}}$ $40.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.41}}$ $60.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.07}}$ $33.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.73}}$
+$\Delta$ $\downarrow$ 2.9% $\downarrow$ 2.8% $\downarrow$ 10.6% $\downarrow$ 3.7% $\downarrow$ 3.8% $\downarrow$ 9.8%
 Avg.
-STONE $28.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ $41.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $34.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.15}}$ $31.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.95}}$ $46.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.60}}$ $36.35{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.23}}$
-w/ ST-TTC $26.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $39.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $24.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ $28.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $43.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.32}}$ $26.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.49}}$
-$\Delta$downarrow 7.6% $\downarrow$ 4.8% $\downarrow$ 28.6% $\downarrow$ 8.0% $\downarrow$ 6.0% $\downarrow$ 26.8%
+STONE $28.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.13}}$ $41.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ $34.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.15}}$ $31.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.95}}$ $46.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.60}}$ $36.35{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.23}}$
+w/ ST-TTC $26.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $39.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $24.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ $28.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.79}}$ $43.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.32}}$ $26.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.49}}$
+$\Delta$ $\downarrow$ 7.6% $\downarrow$ 4.8% $\downarrow$ 28.6% $\downarrow$ 8.0% $\downarrow$ 6.0% $\downarrow$ 26.8%
 E.2 Visualization Case
-We provide more visualization examples of test set predictions to illustrate the effectiveness of our calibration, as shown in Figure 9
-Figure 9: Show case of improvement in spatio-temporal forecasting through our ST-TTC .
+We provide more visualization examples of test set predictions to illustrate the effectiveness of our calibration, as shown in Figure F.1
 Appendix F More Discussion
-F.1 Limitation
+F.1 Distinction between Spatio-Temporal Forecasting and Long-Term Time Series
+As stated in the paper, our work adheres to the common settings of previous short-term and long-term spatio-temporal forecasting studies (e.g., $12\rightarrow 12$, $24\rightarrow 24$) (Shao et al., 2022a, 2024b).
+However, we are fully aware of the $96\rightarrow 96-720$ settings prevalent in current long-term time series forecasting.
+We wish to share our insights regarding this:
+❶There has been a long-standing debate concerning the long-term predictability of time series (Bergmeir, 2024; Brigato et al., 2025), which we do not intend to overly critique here.
+Nevertheless, it is important to note that in spatio-temporal forecasting, specifically in traffic flow theory research, previous studies (Zheng et al., 2025) utilizing residual analysis have indicated that, after minimizing periodicity, the correlation between traffic data beyond one hour and the past hour’s observations is significantly limited for most sensors.
+Furthermore, typical traffic sensor data collection frequency is approximately 5 minutes.
+Therefore, for practical traffic forecasting and decision-making scenarios, which usually focus on the next 1-2 hours (i.e., max 24 steps), our settings are more aligned with real-world applications.
+❷Given that spatio-temporal forecasting can be considered a complex extension of time series forecasting, the additional dimension of sensor count (up to hundreds or thousands) results in an order of magnitude higher data training cost.
+This is a secondary reason why existing spatio-temporal forecasting often considers 12-step settings.
+❸Another notable finding is that a recent study on the accuracy law of deep time series forecasting (Wang et al., 2025b), emerging subsequent to this paper, identifies a significant exponential relationship between the minimum prediction error of current deep forecasting models and the complexity of window series patterns.
+Adopting a Spectral-domain perspective similar to that of this paper, it defines the complexity of series patterns as the total variance of the amplitude spectrum distribution, thereby characterizing the intrinsic heterogeneity of series variations within each relevant window.
+This approach aligns with ours and provides a novel perspective to explain the learnability and effectiveness of the single-step gradient descent in our calibrator.
+While the study focuses solely on univariate time series forecasting, this limitation does not inherently undermine its relevance.
+It is worth noting that the study concludes that mainstream time series models have not yet reached saturation on traffic scenario-based benchmarks.
+However, we hold a different view, arguing that this conclusion primarily stems from the study’s exclusive use of time series forecasting models, while neglecting mainstream spatio-temporal dependency modeling models.
+For further in-depth discussions, we suggest that future research should address this aspect.
+Figure F.1: Show case of improvement in spatio-temporal forecasting through our ST-TTC .
+F.2 Limitation
 In this paper, we propose a novel paradigm for spatio-temporal forecasting: test-time computing.
-We systematically investigate the effective objectives for test-time computation and how to implement efficient computation.
-Based on detailed analysis, we posit that the key to the former lies in designing a lightweight, spatially-aware calibrator, while the key to the latter lies in designing an effective, non-information-leaking calibration parameter learning mechanism.
-Based on these insights, we present our corresponding solution, ST-TTC .
-We first propose a spectral domain calibrator, which appends a lightweight correction module operating in the frequency domain after the backbone network.
-This module calibrates errors by learning small, node-specific amplitude and phase correction factors.
-Furthermore, coupled with a flash gradient update mechanism featuring a streaming memory queue, it ensures universal, rapid, and resource-efficient test-time computing.
 Although there are still many potential areas for improvement, given the superiority and generality of our ST-TTC , we believe this provides a pathway for future exploration of larger-scale and more effective test-time computation.
@@
 The fundamental reason behind this is our belief that true spatio-temporal foundation models do not yet exist.
-Although some preliminary exploratory work has been done [89, 90, 40, 39, 88], they are far from achieving true zero-shot generalization.
+Although some preliminary exploratory work has been done (Yuan et al., 2024a, b; Li et al., 2024a, c; Yu et al., 2024), they are far from achieving true zero-shot generalization.
 However, considering their future inevitability, we believe that further improving the paradigm of test-time computation, especially how to activate and scale the internal capabilities of spatio-temporal foundation models during testing, goes beyond the design philosophy of our proposed learning with calibration.
@@
 However, considering our current efficiency is already sufficiently good, further exploration is needed on how to potentially slow down the test-time computing process to make it more scalable and improve forecasting effectiveness.
-F.2 Future Work
+F.3 Future Work
 Building upon the research direction presented in this paper, we envision future work encompassing two main aspects: